### v1

# Goal-Oriented Communications for Real-time Inference with Two-Way Delay

Abstract
We design a goal-oriented communication strategy for remote inference, where an intelligent model (e.g., a pre-trained neural network) at the receiver side predicts the real-time value of a target signal based on data packets transmitted from a remote location. The inference error depends on both the Age of Information (AoI) and the length of the data packets. Previous formulations of this problem either assumed IID transmission delays with immediate feedback or focused only on monotonic relations where inference performance degrades as the input data ages. In contrast, we consider a possibly non-monotonic relationship between the inference error and AoI. We show how to minimize the expected time-average inference error under two-way delay, where the delay process can have memory. Simulation results highlight the significant benefits of adopting such a goal-oriented communication strategy for remote inference, especially under highly variable delay scenarios.
Index Terms:
Remote Inference, Goal-oriented Communication, Delay with Memory, Markovian Delay, Two-way Delay, Age of Information, and Dynamic Programming.
I Introduction
Intelligent models, e.g. neural networks, have become powerful tools for solving complex computational problems, thanks to advancements in artificial intelligence. These models have not only increased the popularity of today’s most up-to-date technologies but have also played a key role in enabling their current potential. Digital twins, industrial robots, and self-driving cars are prominent examples of such technologies.
An important application of digital twins over long-distance network connections is the monitoring of spacecraft. Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [2, 3]. Digital twins have been used in spacecraft design and operations, ranging from assisting ground control in saving the Apollo 13 astronauts to testing thermal protection on the Space Shuttle.
In recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring. One example is the digital twin implementation for the METU high voltage testing laboratory. [^1] A visual of this implementation, showing a screenshot of the digital twin of the METU high voltage testing laboratory on the left, and a photograph of the actual site on the right, can be accessed at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660. The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
Such systems, involving real-time monitoring, training, etc., often require computation that relies on data sent from remote locations to maintain good performance. Motivated by this, we focus on goal-oriented communication design for the remote inference problem [4, 5]. Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks [6]. This concept emphasizes that networks should focus on the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source. Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8]. This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state). Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [9, 10, 11, 12, 13, 14].
Since we cannot expect the link layer to operate jointly with the application layer for all possible applications, a surrogate metric that the link layer can utilize when deciding which samples to select and transmit is useful. The Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet. We utilize the AoI as an intermediate metric that can be monitored by the link layer, independent of the specific application. A mapping of application performance in relation to the AoI is all that is required for the link layer to function in a goal-oriented manner. This work builds upon the findings presented in the series of papers [4, 16, 5, 17, 1], which first demonstrated the effectiveness of the AoI metric as a surrogate for enhancing communication in the remote tracking or inference of a process.
In this paper, we study a remote inference problem, where an intelligent model infers the real-time value of a target signal based on data transmitted from a remote location. We assume that the transmitter regularly samples a source and forms data packets containing a number of data samples, which we refer to as the packet length, before each transmission. The intelligent model on the receiver side performs inference using these data packets. The inference error is a function of the AoI and the length of the data packet used by the intelligent model. Recent studies [4, 5] have shown that, in various remote inference problems, the inference error (4) does not always increase monotonically with AoI. This contradicts the conventional assumption that a packet with AoI $=0$ necessarily yields better performance than one with AoI $>0$. Therefore, rather than the generate-at-will model [18, 19], we adopt the “selection-from-buffer” model proposed in [4, 5], which enables the scheduler to choose either fresh or stale data samples from the buffer. Moreover, the inference error decreases monotonically as the packet length increases [17]. This improvement comes at the cost of higher resource usage and increased delay. In our communication design, we explore both the possibly non-monotonic relationship between inference error and AoI, as well as the interplay between packet length and transmission delay.
Due to the need for extreme connectivity, next-generation communication networks are expected to involve numerous connections, leading to the existence of multiple routes between any nodes in the network. The relays in such networks may enforce specific routes for certain data flows, which can dynamically change over time to meet the varying demands of multiple flows. For instance, consider a network involving both terrestrial and non-terrestrial connections. A flow can be fully serviced by non-terrestrial connections, terrestrial connections, or a combination of both, depending on conditions such as the priority of the flow, network congestion, or the availability of satellites. Another example is busy industrial environments, where communication routes may be intermittently unavailable due to shadowing from moving objects, causing flows to be rerouted over time. Consequently, goal-oriented communication design for remote inference must be adaptable to delay conditions significantly varying with memory.
To that end, the technical contributions of this paper are:
  • Extending the system models in [5, 17], we formulate and solve a learning and communication co-design problem for remote inference under two-way delay that varies significantly with memory. We consider both time-invariant and time-variable packet length selection scenarios to address the varying computational capabilities of practical systems. The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference system, thereby making the communication goal-oriented.
  • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost semi-Markov decision process (SMDP). Such problems are often solved using dynamic programming [20, 21] and typically do not have closed-form solutions. However, we derive a closed-form solution for this scenario, presented in Theorem 1.
  • We also model the learning and communication co-design problem under time-variable packet length selection as an infinite-horizon average-cost SMDP. We then formulate a Bellman optimality equation and derive a structural result regarding the optimal solution. Motivated by this result, we present a simplified version of the Bellman optimality equation in Theorem 2. Solving the simplified Bellman equation using dynamic programming has significantly lower time complexity compared to solving the original Bellman optimality equation.
  • We conduct two experiments to evaluate the performance of the derived optimal scheduling policies: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which offers a trace-driven evaluation. Simulation results indicate the substantial benefits of adopting a goal-oriented communication strategy for remote inference.
I-A Related Work
The concept of Age of Information (AoI) has attracted significant research interest; see, e.g., [22, 15, 19, 18, 23, 24, 25, 26, 27, 28, 29, 30, 31, 17, 32, 5, 4, 33, 34] and a recent survey [35]. Initially, research efforts were centered on analyzing and optimizing the average AoI and peak AoI in communication networks [15, 19, 18, 23]. Recent research endeavors have revealed that the performance of real-time applications can be modeled as non-linear functions of AoI, leading to the study of optimizing these non-linear functions in control system scenarios [27, 36], remote estimation [16, 25, 30], and remote inference [22, 4, 17, 5]. While a number of studies have analyzed AoI in queuing models, closest to the spirit of this paper is the control of AoI via replacement of exogenous data arrivals with the generation of data ”at will” [37, 19, 38, 39, 24, 26, 27, 28]. A generalization of this approach is to incorporate jointly optimal sampling and scheduling policies to control not only AoI but a more sophisticated end-to-end distortion criterion by using AoI as an auxiliary parameter [16, 25, 33, 17, 4, 5]. While strikingly more demanding of analysis, these formulations take us closer to goal-oriented communication design.
Almost all previous studies on the “generate-at-will” model adopted an assumption that the penalty of information aging is a non-decreasing function of the AoI [16, 25, 19, 38, 39, 24, 26, 27, 28, 34, 33]. However, it was shown in [22, 4, 17, 32, 5] that the monotonicity of information aging depends heavily on the divergence of the time-series data from being a Markov chain. If the input and target data sequences in a system can be closely approximated as a Markov chain, then the penalty increases as the AoI grows; otherwise, if the data sequence significantly deviates from a Markovian structure, the monotonicity assumption does not hold. Following the approach in [22, 4, 17, 32, 5], this paper models the inference error as a possibly non-monotonic function of AoI.
The works most closely related to this paper are [5, 17], which developed scheduling policies for remote inference, considering a possibly non-monotonic dependency between AoI and practical performance. While [5] developed scheduling policies for fixed packet length, the paper [17] jointly optimized packet length and scheduling strategies. Both the paper [5, 17] considered random IID delay for packet transmissions from the transmitter to the receiver and assumed delay-free feedback from the receiver to the transmitter. In practical systems, a non-zero feedback delay exists [40, 41], which affects the performance of the remote inference algorithm. Moreover, due to the envisioned structure of next-generation communication networks, which contain multiple routes between any pair of nodes, the IID delay may not be a good model. This necessitates a model that captures the memory in the delay process, as in [19]. Extending the problem formulations in [5, 17], this paper designs a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay significantly varying with memory.
II System Model
We consider a remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig. 1. On the transmitter side, the source signal $V_{t}$ is regularly sampled from the data source, and the samples are stored in a buffer containing the most recent $B$ samples. The scheduler determines when to transmit and which samples in the buffer to transmit. On the receiver side, a predictor, e.g., a pre-trained neural network, infers the real-time value of a target signal $Y_{t}$ based on the most recently received samples. Such remote inference systems are crucial for many real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and Cyber-Physical Systems (CPS), where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
Figure 1: A remote inference system, where the packet $X_{t-b}^{l}$ represents the group of samples $\big{(}V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1}\big{)}$.
We assume that the system is time-slotted, and it starts operation at time slot $t=0$. We adopt the selection-from-buffer model proposed in [5], where, at each time slot $t$, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer while the oldest sample is discarded. The buffer, with size $B$, is assumed to be initially full, storing the samples $\big{(}V_{0},V_{-1},\ldots,V_{-B+1}\big{)}$. Thus, at any time slot $t$, the buffer contains the most recent $B$ samples, $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$. Before the $i$-th transmission, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{t-b_{i}}^{l_{i}}=\big{(}V_{t-b_{i}},V_{t-b_{i}-1},\ldots,V_{t-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l}$. Here, $l_{i}\in\{1,\ldots,B\}$ and $b_{i}\in\{0,\ldots,B-l_{i}\}$ represents the packet length and buffer position of the $i$-th packet $X_{t-b_{i}}^{l_{i}}$, respectively. Fig. 2 indicates a sample evolution for two consecutive transmission periods. The packet $X_{t-b_{i}}^{l_{i}}$ is submitted at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$. Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$. We assume that the transmitter always waits for feedback before submitting a new packet. In other words, the transmitter remains silent during the time slots between $S_{i}$ and $A_{i}$ for all $i$.
II-A Communication Network Model
The remote inference system operates on top of a communication network with multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver. Each packet submitted into the network is directed to a route based on the network status regarding availability and congestion. If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route. Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length. The delay is at least one time slot for each packet or ACK transmission.
Let the $i$-th epoch consist of the time slots in the interval $[A_{i-1},A_{i})$. Define $C_{i}$ as the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization. The delay state $C_{i}$ represents the network route used in the $i$-th epoch and specifies the transmission and feedback delay distributions. We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$. The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise. Let $T_{i}(l_{i})=D_{i}-S_{i}$ and $F_{i}=A_{i}-D_{i}$ represent the transmission and feedback delays incurred in the $i$-th epoch, respectively.
The receiver detects the delay state $c_{i}$ at each delivery time $D_{i}$ and embeds this information into the ACK message. The ACK notifies the scheduler that the previous packet transmission is complete and that the next packet can be submitted to the network. The network is reliable, meaning no packet or message is lost during transmission.
II-B Inference Error
The Age of Information (AoI) on the receiver side, $\Delta(t)$, is the time difference between the current time $t$ and the generation time $t-\Delta(t)$ of the freshest sample $V_{t-\Delta(t)}$ in the most recently delivered packet $X_{t-\Delta(t)}^{l}$. The AoI at time slot $t$ is determined by
$\Delta(t)=t-S_{i}+b_{i},\text{ if }D_{i}\leq t<D_{i+1}.$ (1)
Figure 2: A sample evolution for two consecutive transmission periods.
Let $d(t)$ represent the length of the most recently delivered packet by time slot $t$, which can be written as follows:
$d(t)=l_{i},\text{ if }D_{i}\leq t<D_{i+1}.$ (2)
We focus on a class of widely used supervised learning algorithms known as Empirical Risk Minimization [42]. We assume that, for each packet length $l\in\{1,2,\ldots,B\}$, the receiver contains a trained neural network $\phi^{*}_{l}\big{(}\cdot\big{)}$ that takes the AoI $\Delta(t)\in\mathbb{Z}^{+}$ and the packet $X_{t-\Delta(t)}^{l}\in\mathcal{V}^{l}$ as inputs and produces an output $a\in\mathcal{A}$ to infer the real-time value of the target signal $Y_{t}\in\mathcal{Y}$. The performance is evaluated through a loss function $L\big{(}\cdot\big{)}:\mathcal{Y}\times\mathcal{A}\mapsto\mathbb{R}$. The incurred loss is $L(y,a)$ when the output $a$ is taken to predict $Y_{t}=y$. Given the AoI $\delta$ and packet length $l$, the training problem to obtain the neural network $\phi^{*}_{l}\big{(}\cdot,\cdot\big{)}$ is formulated as:
$\min_{\phi\big{(}\cdot,\cdot\big{)}\in\Lambda}\mathbb{E}_{Y,X^{l}\sim P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}}\big{[}L\big{(}Y,\phi\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (3)
where $\Lambda$ is the set of all mappings that the neural network can generate, and $P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}$ represents the empirical distribution of the target signal $\tilde{Y}_{0}$ and the packet $\tilde{X}_{-\delta}^{l}$. The AoI $\delta$ is the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}$.
These trained predictors are used to estimate the real-time value of the target signal $Y_{t}$ on the receiver side. We assume that the process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ is stationary, and that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are independent. The average inference error at time slot $t$, given $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
$\varepsilon\big{(}\delta,l\big{)}=\mathbb{E}_{Y,X^{l}\sim P_{Y_{t},X_{t-\delta}^{l}}}\big{[}L\big{(}Y,\phi^{*}_{l}\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (4)
where $P_{Y_{t},X_{t-\delta}^{l}}$ represents the joint distribution of the target signal $Y_{t}$ and the packet $X_{t-\delta}^{l}$.
We are interested in loss functions $L\big{(}\cdot\big{)}$ for which the inference error (4) is bounded. Apart from this restriction, the loss function $L\big{(}\cdot\big{)}$ can be chosen based on the goal of the remote inference system. For example, a quadratic loss function $L_{2}\big{(}y,\hat{y}\big{)}=\lVert y-\hat{y}\rVert_{2}^{2}$ is used in neural network-based minimum mean-squared estimation, where the action $a=\hat{y}$ is an estimate of the target signal $Y_{t}=y$ and $\lVert y\rVert_{2}$ is the Euclidean norm. In softmax regression (i.e., neural network-based maximum likelihood classification), the action $a=\mathcal{Q}_{Y}$ is a distribution of $Y_{t}$, and the loss function $L_{\log}\big{(}y,\mathcal{Q}_{Y}\big{)}=-\log\mathcal{Q}_{Y}(y)$ is the negative log-likelihood function of the value $Y_{t}=y$.
In the subsequent sections, we solve a learning and communication co-design problem that aims to optimize the performance of a remote inference system. From equation (4), we know that the inference error is a function of both the AoI and packet length. For a given AoI $\delta$, the inference error is a non-increasing function of packet length [17]. However, for a given packet length $l$, the inference error does not always increase monotonically with the AoI [4, 5]. This observation contradicts the conventional assumption that fresher packets yield smaller inference errors. Theoretical and experimental explanations of this counter-intuitive observation was provided in [4, 5]. Therefore, we do not adopt the generate-at-will model [18, 19] in this paper.
II-C Structure of a Scheduling Policy
Upon the ACK reception time $A_{i}$, the scheduler determines a waiting time $\tau_{i+1}$ to specify the next submission time slot $S_{i+1}=A_{i}+\tau_{i+1}$. Then, a data packet $X_{S_{i+1}-b_{i+1}}^{l_{i+1}}$ is formed by selecting the packet length $l_{i+1}$ and buffer position $b_{i+1}$. A scheduling policy $\pi\in\Pi$ is defined as a tuple $\pi=\big{(}f,\ell,g\big{)}$, where $f=\big{(}b_{2},b_{3},\ldots\big{)}$ is the buffer position sequence, $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$ is the packet length sequence, $g=\big{(}\tau_{2},\tau_{3},\ldots\big{)}$ is the waiting time sequence, and $\Pi$ denotes the set of all causal policies $\pi$. We assume that the scheduling policy does not employ the knowledge of the observed process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$. This assumption implies that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are independent. The initial conditions of the system are assumed to be $S_{1}=0$, $l_{1}=1$, $b_{1}=0$, $c_{1}\in\{1,2,\ldots,N\}$, and $\Delta(0)$ is a finite constant.
III Learning and Communications Co-design: Time-invariant Packet Length Selection
While designing goal-oriented communication protocols for remote inference, selecting a constant packet length is a simple and often effective choice. If packet lengths vary dynamically from one packet to the next, the receiver needs to reconfigure the estimator over time to accommodate the changing input dimensions, which adds complexity. In this section, we consider systems where this complexity is avoided by choosing a time-invariant packet length sequence. Systems that can vary the packet length over time are discussed in Section IV.
III-A Co-design Problem Formulation
Let $\Pi_{l}$ represent the set of all causal scheduling policies $\pi=(f,\ell,g)$ with a time-invariant packet length sequence =${\ell=\{l,l,\ldots\}}$ for $l\in\{1,2,\ldots,B\}$:
$\Pi_{l}=\{\pi\in\Pi:l=l_{2}=l_{3}=\ldots\}.$ (5)
The learning and communications co-design problem with time-invariant packet length selection can be formulated as a two-layer nested optimization problem:
$\varepsilon_{l,\text{opt}}=\inf_{\pi\in\Pi_{l}}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{T-1}\varepsilon\big{(}\Delta(t),l\big{)}\Bigg{]},$ (6)
$\varepsilon_{\text{opt}}^{\text{inv.}}=\min_{l\in\{1,2,\ldots,B\}}\varepsilon_{l,\text{opt}},$ (7)
where $\varepsilon_{l,\text{opt}}$ is the optimum value of the inner optimization problem (6) for a given packet length $l$, and $\varepsilon_{\text{opt}}^{\text{inv.}}$ is the optimum value of the two-layer nested optimization problem (6)-(7) with the optimal time-invariant packet length.
The inner optimization problem (6) can be cast as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP). While dynamic programming algorithms, namely policy iteration and value iteration, are typically employed to solve such problems [20, 21], we are able derive a closed-form solution that achieves significantly lower computational complexity than dynamic programming. Moreover, the outer optimization problem (7) is a straightforward search over the integer values $\{1,2,\ldots,B\}$.
Unlike the formulation in [17], this problem formulation accounts for varying delay conditions between the transmitter and the receiver and does not assume immediate feedback.
III-B Optimal Solution to (6)-(7)
We define each ACK reception time $A_{i}$ as a decision time for the SMDP (6). The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch. The actions are the waiting time $\tau_{i+1}$ and the buffer position $b_{i+1}$. A detailed description of the SMDP can be found in Appendix A. The Bellman optimality equation of this SMDP can be formulated as:
$h_{1}\big{(}\delta,c_{i}\big{)}=\min_{\begin{subarray}{c}\tau_{i+1}\in\{0,1,\ldots\}\\ b_{i+1}\in\{0,1,\ldots,B-l\}\end{subarray}}$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\big{[}h_{1}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]},$ (8)
where $h_{1}(\delta,c_{i})$ is the relative value function. We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6). This involves determining the optimal buffer position $b_{i+1}^{*}$ and waiting time $\tau_{i+1}^{*}$ for each state $\big{(}\delta,c_{i}\big{)}$. By exploiting the structural properties of the Bellman optimality equation (III-B), we are able to obtain a closed-form solution to (6), as asserted in the following theorem.
Define the index function
$\gamma\big{(}\delta,d,l,c\big{)}=\min_{\tau\in\{1,2,\ldots\}}\frac{1}{\tau}\sum_{k=0}^{\tau-1}\mathbb{E}\bigg{[}\varepsilon\big{(}\delta+T_{i+1}(l)+k,d\big{)}\bigg{|}C_{i}=c\bigg{]},$ (9)
for all $\delta\in\{1,2,\ldots\}$, $d,l\in\{1,2,\cdot,B\}$, and $c\in\{1,2,\ldots,N\}$.
Theorem 1.
There exists an optimal solution $\pi^{*}=\big{(}\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)},\big{(}l,l,\ldots\big{)},\big{(}\tau_{2}^{*},\tau_{3}^{*},\ldots\big{)}\big{)}\in\Pi_{l}$ to problem (6), where the optimal waiting time $\tau_{i+1}^{*}$ is determined by the index-based threshold rule
$\tau_{l}\big{(}\delta,c_{i}\big{)}=\min\{k\geq 0:\gamma\big{(}\delta+k,l,l,c_{i}\big{)}\geq\beta\},$ (10)
the optimal buffer position $b_{i+1}^{*}$ is given by
$b_{i+1}^{*}=\operatorname*{arg\,min}_{b\in\{0,1,\ldots,B-l\}}$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\mathfrak{D}-1}\bigg{(}\varepsilon\big{(}b+T_{i+1}(l)+k,l\big{)}-\beta\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]},$ (11)
$\mathfrak{D}=F_{i+1}+\tau_{l}\big{(}b+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, and $\gamma\big{(}\cdot\big{)}$ is defined in (9). Furthermore, $\beta$ is the unique root of
$\mathbb{E}\Bigg{[}\sum_{t=A_{i}(\beta)}^{A_{i+1}(\beta)-1}\varepsilon\big{(}\Delta(t)\big{)}\Bigg{]}-\beta\mathbb{E}\big{[}A_{i+1}(\beta)-A_{i}(\beta)\big{]}=0,$ (12)
and it is exactly the optimum value $\varepsilon_{l,\text{opt}}$ of problem (6).
Proof Sketch.
We prove Theorem 1 in three steps:
Step 1: We obtain two key results by using the Bellman optimality equation (III-B):
  • The optimal waiting time $\tau_{i+1}^{*}$ is determined by the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ given by equation (10) for all $i$.
  • The delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining the optimal buffer position $b_{i+1}^{*}$ for all $i$.
Step 2: Since the delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining $b_{i+1}^{*}$ for all $i$, we can construct a new SMDP for determining $b_{i+1}^{*}$ using $c_{i}$ as the state. In the new SMDP, we fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. Moreover, in the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. The Bellman optimality equation of this new SMDP is expressed as:
$h_{1}^{\prime}\big{(}c_{i}\big{)}=\min_{\begin{subarray}{c}b_{i+1}\in\{0,1,\ldots,B-l\}\end{subarray}}$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\vartheta_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\bigg{[}h_{1}^{\prime}\big{(}C_{i+1}\big{)}\bigg{|}C_{i}=c_{i}\bigg{]},$ (13)
where $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, and $h_{1}^{\prime}\big{(}c_{i}\big{)}$ is the relative value function.
The Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision-epoch optimization problem, allowing any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ to be expressed as shown in (1).
Step 3: We determine that the optimal value $\varepsilon_{l,\text{opt}}$ for the inner optimization problem (6) is the unique root of (12).
The detailed proof is provided in Appendix B. ∎
The optimal scheduling policy for the inner optimization problem (6), outlined in Theorem 1, is well-structured. Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. The index function $\gamma\big{(}\cdot\big{)}$, provided in (9), can be readily calculated for any state $\big{(}\delta,c_{i}\big{)}$, using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$.
Furthermore, according to equation (1), $b_{i+1}^{*}$ is designed to minimize the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10). Also, each optimal buffer position $b_{i+1}^{*}$ depends solely on the delay state $c_{i}$ in the $i$-th epoch, and is independent of $\delta=\Delta(A_{i})$. This result can be better understood by deeply analyzing the AoI evolution. At each delivery time slot $D_{i+1}$, the AoI resets to a realization of $b_{i+1}^{*}+T_{i+1}(l)$. Due to these regular resets, the AoI evolution between consecutive delivery slots $D_{i+1}$ and $D_{i+2}$ is independent of the AoI evolution outside this interval. Moreover, $b_{i+1}^{*}$ is a parameter to control the AoI evolution between $D_{i+1}$ and $D_{i+2}$.
Finally, the optimum value $\varepsilon_{l,\text{opt}}$, which is the threshold in the waiting time rule (10) and necessary for solving (1), is the unique root of equation (12). This equation can be efficiently solved using low-complexity algorithms [25, Algorithms 1-3], such as bisection search.
The optimal policy in Theorem 1 has two key distinctions compared to the policy established in [17, Theorem 1] for remote inference under time-invariant packet length selection, IID transmission delays, and immediate feedback. First, although the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ in (10) closely resembles the waiting time rule derived in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state. For the problem formulation (6)-(7), which involves a finite number of delay states with non-zero feedback delay, the existence of such a waiting time rule was not evident and required additional technical efforts for its derivation.
Second, in [17, Theorem 1], the optimal buffer position sequence is time-invariant due to the assumption of IID transmission delays. However, when the delay state changes with memory, achieving optimal performance requires a buffer position sequence that can adapt to variations in delay statistics. Theorem 1 unveils that the optimal buffer position sequence is determined by solving $N$ per-sample optimization problems, each related to a distinct delay state, without requiring pre-computation of the relative value function $h_{1}\big{(}\cdot\big{)}$.
We can solve the inner optimization problem (6) for any given packet length $l$ by using Theorem 1. Then, the outer optimization problem (7) is just a straightforward search over the integer values $\{1,2,\ldots,B\}$, and $\varepsilon_{\text{opt}}^{\text{inv.}}=\varepsilon_{l^{*},\text{opt}}$.
IV Learning and Communications Co-design: Time-Variable Packet Length Selection
In contrast to the previous case, where high computational complexity is avoided, here we aim to achieve optimal performance without any restrictions on the packet length sequence $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$. This unconstrained approach allows us to maximize the system’s potential by dynamically adjusting the packet length $l_{i}$ over time. Therefore, this case represents the most general solution.
IV-A Co-design Problem Formulation
The learning and communications co-design problem with time-variable packet length selection is formulated as:
$\varepsilon_{\text{opt}}=\inf_{\pi\in\Pi}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{T-1}\varepsilon\big{(}\Delta(t),d(t)\big{)}\Bigg{]},$ (14)
where $\varepsilon_{\text{opt}}$ is the optimum value of problem (14).
Problem (14) is an infinite-horizon average-cost SMDP, which is more challenging to solve compared to problem (6). We formulate a Bellman optimality equation for this SMDP and then simplify it using a structural result regarding the optimal policy. The time complexity for solving the simplified Bellman optimality equation using dynamic programming algorithms is substantially lower than that for solving the original Bellman optimality equation.
IV-B Optimal Solution to (14)
Each ACK reception time $A_{i}$ is a decision time for the SMDP (14). The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,d,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch. The actions to be taken at each decision time $A_{i}$ are the waiting time $\tau_{i+1}$, the packet length $l_{i+1}$, and the buffer position $b_{i+1}$. A detailed description of the SMDP can be found in Appendix C. The Bellman optimality equation of this SMDP can be expressed as:
$h_{2}\big{(}\delta,d,c_{i}\big{)}=\min_{\begin{subarray}{c}\tau_{i+1}\in\{0,1,\ldots\}\\ l_{i+1}\in\{1,2,\ldots,B\}\\ b_{i+1}\in\{0,1,\ldots,B-l_{i+1}\}\end{subarray}}$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l_{i+1})-1}\bigg{(}\varepsilon\big{(}\delta+k,d\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]},$ (15)
where $h_{2}\big{(}\delta,d,c_{i}\big{)}$ is the relative value function.
By solving the Bellman optimality equation (IV-B) using dynamic programming, we can obtain a solution to problem (14). The time complexity of this dynamic programming algorithm is high, as it requires the joint optimization of three variables for each state $\big{(}\delta,d,c_{i}\big{)}$. Theorem 2 provides a threshold rule for determining the waiting time action at each decision time $A_{i}$ and simplifies the Bellman optimality equation. The time complexity for solving the simplified Bellman optimality equation is significantly lower than that for solving (IV-B).
Theorem 2.
There exists an optimal solution $\pi^{*}=\big{(}\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)},\big{(}l_{2}^{*},l_{3}^{*},\ldots\big{)},\big{(}\tau_{2}^{*},\tau_{3}^{*},\ldots\big{)}\big{)}\in\Pi$ to problem (14). The optimal waiting time $\tau_{i+1}^{*}$ is determined by the index-based threshold rule
$\tau\big{(}\delta,d,l_{i+1},c_{i}\big{)}=\min\{k\geq 0:\gamma\big{(}\delta+k,d,l_{i+1},c_{i}\big{)}\geq\varepsilon_{\text{opt}}\},$ (16)
where $\gamma\big{(}\cdot\big{)}$ is defined in (9). Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the simplified Bellman optimality equation
[…]
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau\big{(}\delta,d,l_{i+1},c_{i}\big{)}+T_{i+1}(l_{i+1})-1}\bigg{(}\varepsilon\big{(}\delta+k,d\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
[…]
$\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]}\big{\}}\big{\}}.$ (17)
Proof Sketch.
Theorem 2 is proved motivated by the observation that the optimal waiting time $\tau_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ for any state $\big{(}\delta,d,c_{i}\big{)}$ can be determined by solving separate optimization problems, given the optimal packet length $l_{i+1}^{*}$. Subsequently, the solution to the waiting time problem is derived as an index-based threshold rule, which allows us to simplify the Bellman optimality equation (IV-B) to (2). The detailed proof is provided in Appendix E. ∎
Both Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [20, Chapter 11.4.4]. Policy iteration is efficient for solving SMDPs [17, 20]. The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B). The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2). The policy evaluation step is identical for both equations.
V Simulations
To evaluate the performance of the optimal scheduling policies outlined in Theorems 1 and 2, we consider two experiments: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which provides a trace-driven evaluation.
V-A Model-Based Evaluation
Control systems with memory and delay [43, 44, 45] and wireless channels [46] are often modeled as higher-order autoregressive (AR) linear time-invariant systems. Motivated by this, we consider a remote inference problem for an AR process, where the target evolves according to:
$Y_{t}=a_{1}Y_{t-1}+a_{2}Y_{t-2}+\ldots+a_{p}Y_{t-p}+W_{t},$ (18)
where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$ and $Y_{t}\in\mathbb{R}$. Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of the target $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$. We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$. The remaining coefficients are zero. We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$. In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$. Since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic, the optimal inference error is achieved by a linear MMSE estimator. Fig. 3 shows the average inference error for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
Figure 3: Inference error of the AR process experiment for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
Using this AR process experiment, we illustrate the performance of the optimal policy derived for the scheduling problem (6)-(7) under time-invariant packet length selection. Let the number of delay states $N$ be $2$. The transmission delay $T_{i}(l)$ corresponding to each state is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (19)
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$. We adjust the delay memory using the parameter $\alpha=p_{12}+p_{21}$, which is equal to the sum of transition probabilities of the Markov chain according to which the delay state evolves. We set $p_{12}=p_{21}$, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in the steady state for any $\alpha$. The maximum allowed packet length is 10, and the buffer size $B=75$.
Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $2$ with $\alpha=\frac{1}{20}$:
  • 1.
    Optimal policy derived for the scheduling problem (6)-(7) under time-invariant packet length selection.
  • 2.
    Policy outlined in Theorem 1 for $l=1$.
  • 3.
    Generate-at-will, zero-wait, $l=1$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(1,1,\ldots)$.
The result in Fig. 4 highlights that the first policy achieves a time-average inference error up to 6 times smaller compared to the third policy. Although part of this gain can be obtained with the packet length $l=1$ by optimally determining the waiting times and buffer positions as in the second policy, a larger portion of the gain comes from selecting the appropriate packet length, accounting for the interplay between the transmission delay statistics and the packet length $l$.
Figure 4: Time-average inference error vs. the scaling parameter $\sigma$ of transmission delay $T_{i}(l)$.
Fig. (5) presents the time-average inference error achieved by the following two scheduling policies as the parameter $\alpha$ varies from 00 to $2$ with $\sigma=\frac{5}{2}$:
  • 1.
    Policy outlined in Theorem 1 for $l=5$.
  • 2.
    Policy outlined in [17, Theorem 1] for a fixed packet length of $l=5$. During the simulations, this policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
The result in Fig. (5) emphasizes the importance of considering delay memory when designing scheduling algorithms. It demonstrates that the first policy offers minimal improvement compared to the second policy when $\alpha$ is close to $1$. This is because the delay distribution becomes exactly IID at $\alpha=1$. However, as $\alpha$ deviates from this value, the delay memory increases, and the advantage of the first policy becomes apparent, with a significant performance gain of up to $11.6\%$.
Figure 5: Time-average inference error vs. the sum of transition probabilities $\alpha=p_{12}+p_{21}$.
V-B Trace-Driven Evaluation
We consider the OpenAI CartPole-v1 task [47], where a Deep Q-Network (DQN) reinforcement learning algorithm [48] is used to control the force applied to a cart to prevent the attached pole from falling over. A pre-trained DQN neural network from [5] is employed for this control task. A time-series dataset containing the pole angle $\psi_{t}$ and cart position $Z_{t}$ is obtained from [5]. This dataset was generated by simulating 10,000 episodes of the OpenAI CartPole-v1 environment using the DQN controller. The objective is to predict the pole angle and position $Y_{t}=(\psi_{t},Z_{t})$ at time $t$ based on the feature vector $X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta+l-1},Z_{t-\delta+l-1})\big{)}$, which represents a sequence of past pole angles and cart positions generated $\delta$ time slots ago, with a sequence length of $l$.
Figure 6: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$.
Ensuring the safe operation of industrial and robotic systems often requires evaluating their state based on observations from remote sources. Consequently, the prediction problem is critical for monitoring the safety of these systems. We focus on a specific scenario where a camera remotely monitors a cart-pole system. The camera captures video frames and transmits them to a receiver. Due to the high dimensionality of the video data, transmitting the sequence of frames requires multiple time slots. Consequently, the most recently delivered sequence of video frames is generated $\delta$ time slots ago. Once received, the frames undergo signal processing to extract relevant information about the cart’s position and the pole’s angle. We assume this signal processing is accurate. Then, a predictor takes the most recent available observations and predicts the current pole angle and cart position $Y_{t}=(\psi_{t},Z_{t})$.
The predictor in this experiment is a Long Short-Term Memory (LSTM) neural network consisting of one input layer, one hidden layer with 64 LSTM cells, and a fully connected output layer. The first 72% of the dataset is used for training, and the remaining data is used for inference. The performance of the trained LSTM predictor is evaluated on the inference dataset. The inference error is shown in Fig. 6 for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$. We observe that the inference error is not a monotonic function of AoI $\delta$ for a given feature vector length $l$. Additionally, the inference error decreases as the feature vector length $l$ increases for a fixed AoI value $\delta$.
Using this cart-pole experiment, we illustrate the performance of the optimal policy outlined in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection. We again consider two delay states, i.e., ${N=2}$. The transmission delay $T_{i}(l)$ is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (20)
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the parameter $\alpha=p_{12}+p_{21}=\frac{1}{20}$. The transition probabilities $p_{12}$ and $p_{21}$ are equal, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in steady state for any $\alpha$.
(a) $\sigma=\frac{1}{2}$
(b) $\sigma=\frac{5}{2}$
Figure 7: Time-average inference error vs. the buffer size B for ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$.
Figure 8: Time-average inference error vs. the scaling parameter $\sigma$ of transmission delay $T_{i}(l)$.
Fig. 7(a) and Fig. 7(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$, respectively. As the buffer size increases, the performance improvement diminishes and becomes less significant. Therefore, we fix the buffer size $B$ at $10$ for this experiment.
Fig. 8 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $3$:
  • 1.
    Optimal policy in Theorem 2
  • 2.
    Generate-at-will, zero-wait, $l=1$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(1,1,\ldots)$.
  • 3.
    Generate-at-will, zero-wait, $l=5$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(5,5,\ldots)$.
The result in Fig. 8 highlights the importance of adjusting the packet length over time in response to varying delay conditions. The optimal policy in Theorem 2 outperforms the other two policies, which use constant packet length, achieving up to $25\%$ reduction in average inference error.
VI Conclusion
This paper studies a goal-oriented communication design problem for remote inference, where an intelligent model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location. We derive two optimal scheduling policies under time-invariant and time-variable packet length selection. These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory. Finally, through the simulations, we demonstrate the significant benefits of adopting such a goal-oriented communication strategy for remote inference.
Acknowledgment
The authors would like to thank Batu Saatci for his assistance with the simulations.
Appendix A
The SMDP corresponding to problem (6), constructed in Section III-B, is described by the following components [21]:
(i) Decision Time: Each ACK reception time slot $A_{i}=S_{i}+T_{i}(l)+F_{i}$ is a decision time.
(ii) Action: At each decision time $A_{i}$, the scheduler determines the waiting time $\tau_{i+1}$ and buffer position $b_{i+1}$ for the next packet to be transmitted.
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big{(}\delta,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ evolves according to equation (1). Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
(iv) Transition Time and Cost: The time between the consecutive decision times $A_{i+1}-A_{i}$ is represented by the random variable $\tau_{i+1}+T_{i+1}(l)+F_{i+1}$, given the delay state $C_{i}=c_{i}$. Moreover, the cost incurred while transitioning from decision time $A_{i}$ to decision time $A_{i+1}$,
$\sum_{t=A_{i}}^{A_{i+1}-1}\varepsilon\big{(}\Delta(t),l\big{)},$
is represented by the random variable
$\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l)-1}\varepsilon\big{(}\Delta(A_{i})+k,l\big{)}+\sum_{k=0}^{F_{i+1}-1}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)},$
given the delay state $C_{i}=c_{i}$.
Appendix B Proof of Theorem 1
This proof consists of three steps.
Step 1: We prove two structural results regarding the optimal scheduling policy by using the Bellman optimality equation (III-B), which comprises three terms. The first term,
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]},$
is independent of the buffer position $b_{i+1}$. In contrast, the remaining two terms,
$\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
and
$\mathbb{E}\bigg{[}h_{1}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}\bigg{|}C_{i}=c_{i}\bigg{]},$
are independent of the waiting time $\tau_{i+1}$. Consequently, the optimal buffer position $b_{i+1}^{*}$ and the optimal waiting time $\tau_{i+1}^{*}$ can be determined by solving separate optimization problems.
(i) Optimal Waiting Time: The optimal waiting time $\tau_{i+1}^{*}$, when $\Delta(A_{i})=\delta$ and $C_{i}=c_{i}$, can be determined by solving the optimization problem
$\min_{\tau_{i+1}\in\{0,1,\ldots\}}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}.$ (21)
Problem (21) is a simple integer optimization problem. The optimal waiting time $\tau_{i+1}^{*}=0$ if
$\min_{\tau_{i+1}\in\{1,2,\ldots\}}$ $\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}\geq$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}.$ (22)
By some algebraic manipulations, we can obtain
$\min_{\tau_{i+1}\in\{1,2,\ldots\}}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}-1}\bigg{(}\varepsilon\big{(}\delta+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$\geq 0.$ (23)
The inequality (B) holds if and only if
$\min_{\tau_{i+1}\in\{1,2,\ldots\}}\frac{1}{\tau_{i+1}}\sum_{k=0}^{\tau_{i+1}-1}\mathbb{E}\bigg{[}\varepsilon\big{(}\delta+T_{i+1}(l)+k,l\big{)}\bigg{|}C_{i}=c_{i}\bigg{]}\geq\varepsilon_{l,\text{opt}}.$ (24)
The left-hand side of (24) is equal to $\gamma\big{(}\delta,l,l,c_{i}\big{)}$, where $\gamma\big{(}\cdot\big{)}$ is the index function defined in (9). Similarly, the optimal waiting time $\tau_{i+1}^{*}=1$, if $\tau_{i+1}^{*}\neq 0$ and
$\min_{\tau_{i+1}\in\{2,3,\ldots\}}\mathbb{E}\Bigg{[}\sum_{k=1}^{\tau_{i+1}-1}\bigg{(}\varepsilon\big{(}\delta+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$\geq 0.$ (25)
Following the same steps in (B)-(24), we can demonstrate that the inequality (B) is equivalent to
$\gamma\big{(}\delta+1,l,l,c_{i}\big{)}\geq\varepsilon_{l,\text{opt}}.$ (26)
By repeating (B)-(26), we can establish the rule that the optimal waiting time $\tau_{i+1}^{*}$ is equal to $k$, provided that ${\tau_{i+1}^{*}\neq 0,1,\ldots,k-1}$ and
$\gamma\big{(}\delta+k,l,l,c_{i}\big{)}\geq\varepsilon_{l,\text{opt}}.$ (27)
Hence, any $\tau_{i+1}^{*}$ in the optimal waiting time sequence ${g^{*}=\big{(}\tau_{2}^{*},\tau_{3}^{*},\ldots\big{)}}$ is determined by the index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ given by equation (10).
(ii) Optimal Buffer Position: The optimal buffer position $b_{i+1}^{*}$, when $\Delta(A_{i})=\delta$ and $C_{i}=c_{i}$, can be determined by solving the optimization problem
$\min_{b_{i+1}\in\{0,1,\ldots,B-l\}}$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
$+\mathbb{E}\bigg{[}h_{1}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}\bigg{|}C_{i}=c_{i}\bigg{]}.$ (28)
From (B), we observe that the optimal buffer position $b_{i+1}^{*}$ is independent of $\delta=\Delta(A_{i})$ and depends only on the delay state $c_{i}$ in the previous epoch. In other words, $c_{i}$ is a sufficient statistic for determining $b_{i+1}^{*}$ for all $i$.
Step 2: We construct a new SMDP based on the results obtained in the first step. In the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. At each decision time $D_{i+1}$, the system state is represented only by the delay state $c_{i}$ in the $i$-th epoch. We fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$, and the action involves determining the buffer position $b_{i+1}$. A detailed description of this SMDP can be found in Appendix D. The Bellman optimality equation of this new SMDP is given by equation (III-B).
The second term on the right-hand side of equation (III-B),
$\mathbb{E}\bigg{[}h_{1}^{\prime}\big{(}C_{i+1}\big{)}\bigg{|}C_{i}=c\bigg{]},$
is independent of the action $b_{i+1}$ taken at decision time $D_{i+1}$ and depends only on the state $C_{i}=c_{i}$. Therefore, the Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision epoch optimization problem; that is, we can express any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ as shown in (1).
Step 3: We determine the optimal value $\varepsilon_{l,\text{opt}}$ of the inner optimization problem (6) as the unique root of equation (12). Let $\vartheta_{i}^{*}$ denote the random variable $b_{i}^{*}+T_{i}(l)+F_{i}$. Consider the following equality:
$\mathbb{E}\Bigg{[}\sum_{t=A_{i}}^{A_{i+1}-1}\varepsilon\big{(}\Delta(t)\big{)}\Bigg{]}-\varepsilon_{l,\text{opt}}\mathbb{E}\big{[}A_{i+1}-A_{i}\big{]}=$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{l}\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}.$ (29)
We first demonstrate that the right-hand side of equation (B) is equal to zero, which implies $\varepsilon_{l,\text{opt}}$ is indeed a root of equation (12). From the Bellman equation (III-B), we can obtain the following relation:
$h_{1}\big{(}\vartheta_{i}^{*},C_{i}\big{)}=$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{l}\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}\Bigg{]}$
$+\mathbb{E}\bigg{[}h_{1}\big{(}\vartheta_{i+1}^{*},C_{i+1}\big{)}\bigg{|}C_{i}\bigg{]}.$ (30)
By the law of iterated expectations, taking the expectation over $C_{i}$ on both sides of (B) gives us the equation
$\mathbb{E}\big{[}h_{1}\big{(}\vartheta_{i}^{*},C_{i}\big{)}\big{]}=$
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{l}\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}$
$+\mathbb{E}\big{[}h_{1}\big{(}\vartheta_{i+1}^{*},C_{i+1}\big{)}\big{]}.$ (31)
Since $C_{i}$ evolves according to an ergodic Markov chain with a unique stationary distribution, it follows that $\mathbb{E}\big{[}h_{1}\big{(}\vartheta_{i}^{*},C_{i}\big{)}\big{]}=\mathbb{E}\big{[}h_{1}\big{(}\vartheta_{i+1}^{*},C_{i+1}\big{)}\big{]}$. Therefore, equation (B) implies that the right-hand side of equation (B) is equal to zero.
Next, we prove the uniqueness of the root $\varepsilon_{l,\text{opt}}$. The right-hand side of equation (B) can be written as
$\min_{\tau\big{(}\cdot\big{)}}\text{ }\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}$
$+\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}.$ (32)
The expression in (B) is concave, continuous, and strictly decreasing in $\varepsilon_{l,\text{opt}}$, as the first term is the functional minimum of linear decreasing functions of $\varepsilon_{l,\text{opt}}$, and the second term is itself a linear decreasing function of $\varepsilon_{l,\text{opt}}$. Therefore, since
$\lim_{\varepsilon_{l,\text{opt}}\to\infty}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}=-\infty$
and
$\lim_{\varepsilon_{l,\text{opt}}\to-\infty}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau\big{(}\vartheta_{i}^{*},C_{i}\big{)}+T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\vartheta_{i}^{*}+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}=\infty$
for any $\tau\big{(}\cdot\big{)}$, and
$\lim_{\varepsilon_{l,\text{opt}}\to\infty}\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}=-\infty$
and
$\lim_{\varepsilon_{l,\text{opt}}\to-\infty}\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}^{*}+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{]}=\infty,$
equation (12) has a unique root. This completes the proof.
Appendix C
The SMDP corresponding to problem (14), constructed in Section IV-B, is described by the following components [21]:
(i) Decision Time: Each ACK reception time slot $A_{i}=S_{i}+T_{i}(l_{i})+F_{i}$ is a decision time.
(ii) Action: At each decision time $A_{i}$, the scheduler determines the waiting time $\tau_{i+1}$, packet length $l_{i+1}$, and buffer position $b_{i+1}$ for the next packet to be transmitted.
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big{(}\delta,d,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ and the length $d(t)$ of the packet most recently delivered by time slot $t$ evolve according to equations (1) and (2), respectively. Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
(iv) Transition Time and Cost: The time between the consecutive decision times $A_{i+1}-A_{i}$ is represented by the random variable $\tau_{i+1}+T_{i+1}(l_{i+1})+F_{i+1}$, given the delay state $C_{i}=c_{i}$. Moreover, the cost incurred while transitioning from decision time $A_{i}$ to decision time $A_{i+1}$,
$\sum_{t=A_{i}}^{A_{i+1}-1}\varepsilon\big{(}\Delta(t),d(t)\big{)},$
is represented by the random variable
$\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l_{i+1})-1}\varepsilon\big{(}\Delta(A_{i})+k,d(A_{i})\big{)}$
$+\sum_{k=0}^{F_{i+1}-1}\varepsilon\big{(}b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big{)},$
given the delay state $C_{i}=c_{i}$.
Appendix D
The components of the SMDP constructed in the second step of the proof of Theorem (1) differ from those of the SMDPs constructed in Sections III-B and IV-B.
(i) Decision Time: Each delivery time slot $D_{i+1}=S_{i+1}+T_{i+1}(l)$ is a decision time.
(ii) Action: At each decision time $D_{i+1}$, the scheduler determines the buffer position $b_{i+1}$ for the next packet to be transmitted.
(iii) State and State Transitions: The system state at each decision time $D_{i+1}$ is the delay state $C_{i}=c_{i}$ in the $i$-th epoch, which evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
(iv) Transition Time and Cost: The time between consecutive decision times, $D_{i+1}-D_{i}$, is represented by the random variable $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, given the delay state $C_{i}=c_{i}$, where $\tau_{l}\big{(}\cdot\big{)}$ is the waiting time rule defined in (10). Moreover, the cost incurred while transitioning from decision time $D_{i}$ to decision time $D_{i+1}$,
$\sum_{t=D_{i}}^{D_{i+1}-1}\varepsilon\big{(}\Delta(t),l\big{)},$
is represented by the random variable
$\sum_{k=0}^{\vartheta_{i+1}-1}\varepsilon\big{(}b_{i+1}+T_{i+1}(l)+k,l\big{)},$
given the delay state $C_{i}=c_{i}$.
Appendix E Proof of Theorem 2
The Bellman optimality equation in (IV-B) has three terms. The first term,
$\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l_{i+1})-1}\bigg{(}\varepsilon\big{(}\delta+k,d\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]},$
does not depend on the buffer position $b_{i+1}$. On the contrary, the remaining two terms,
$\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
and
$\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]},$
do not depend on the waiting time $\tau_{i+1}$. Therefore, given the optimal packet length $l_{i+1}^{*}$, the optimal waiting time $\tau_{i+1}^{*}$, when $\Delta(A_{i})=\delta$, $d(A_{i})=d$, and $C_{i}=c_{i}$, can be determined by solving the following optimization problem, which is independent of the action $b_{i+1}$:
$\min_{\tau_{i+1}\in\{0,1,\ldots\}}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}+T_{i+1}(l_{i+1}^{*})-1}\bigg{(}\varepsilon\big{(}\delta+k,d\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}.$ (33)
Problem (33) is a simple integer optimization problem, quite similar to problem (21). By following the steps in (B) through (26), we can show that the optimal waiting time $\tau_{i+1}^{*}=k$, if $\tau_{i+1}^{*}\neq 0,1,\ldots,k-1$ and
$\gamma(\delta+k,d,l_{i+1}^{*},c_{i})\geq\varepsilon_{\text{opt}},$ (34)
where $\gamma\big{(}\cdot\big{)}$ is the index function defined in (9). Hence, the optimal waiting time $\tau_{i+1}^{*}$ is determined as given in equation (16), and the Bellman optimality equation (IV-B) can be equivalently formulated as (2). This completes the proof.

### v3

@@
-Goal-Oriented Communications for Real-time Inference with Two-Way Delay
+Goal-Oriented Status Updating for Real-time Remote Inference over Networks with Two-Way Delay
 Abstract
-We design a goal-oriented communication strategy for remote inference, where an intelligent model (e.g., a pre-trained neural network) at the receiver side predicts the real-time value of a target signal based on data packets transmitted from a remote location.
-The inference error depends on both the Age of Information (AoI) and the length of the data packets.
-Previous formulations of this problem either assumed IID transmission delays with immediate feedback or focused only on monotonic relations where inference performance degrades as the input data ages.
-In contrast, we consider a possibly non-monotonic relationship between the inference error and AoI.
-We show how to minimize the expected time-average inference error under two-way delay, where the delay process can have memory.
-Simulation results highlight the significant benefits of adopting such a goal-oriented communication strategy for remote inference, especially under highly variable delay scenarios.
+We study a setting where an intelligent model (e.g., a pre-trained neural network) predicts the real-time value of a target signal using data samples transmitted from a remote source according to a scheduling policy.
+The scheduler decides on i) the age of the samples to be sent, ii) when to send them, and iii) the length of each packet (i.e., the number of samples contained in each packet).
+The dependence of inference quality on the Age of Information (AoI) for a given packet length is modeled by a general relationship.
+Previous work assumed i.i.d.
+transmission delays with immediate feedback or were restricted to the case where inference performance degrades as the input data ages.
+Our formulation, in addition to capturing non-monotone age dependence, also covers Markovian delay on both forward and feedback links.
+We model this as an infinite-horizon average-cost Semi-Markov Decision Process.
+We obtain a closed-form solution that decides on (i) and (ii) for any constant packet length.
+The solution for when to send is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver.
+The age of the packet selected is a function of the delay state.
+We separately optimize the value of the constant length.
+We also develop an index-based threshold policy for the variable length case, which allows a complexity reduction.
+In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
 Index Terms:
@@
 I Introduction
-Intelligent models, e.g.
-neural networks, have become powerful tools for solving complex computational problems, thanks to advancements in artificial intelligence.
-These models have not only increased the popularity of today’s most up-to-date technologies but have also played a key role in enabling their current potential.
-Digital twins, industrial robots, and self-driving cars are prominent examples of such technologies.
-An important application of digital twins over long-distance network connections is the monitoring of spacecraft.
+Today, many previously complex computational problems enjoy the aid of AI.
+Among these AI-aided approaches are intelligent models, e.g., neural networks.
+These models have played a key role in enabling certain technologies as digital twins, industrial robotics, and self-driving cars.
+However, these models rely on data to be delivered to them, which in the remote computation context, depends on the network delivering the relevant data in a timely fashion.
+This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2].
+An extreme example of monitoring over long-distance network connections is a spacecraft digital twin.
 Consider training a digital twin on Earth based on data coming from rovers on Mars or calibrating the virtual replica of the spacecraft/satellite using live telemetry data [3, 4].
-Digital twins have been used in spacecraft design and operations, ranging from assisting ground control in saving the Apollo 13 astronauts to testing thermal protection on the Space Shuttle.
-In recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring.
-One example is the digital twin implementation for the METU high voltage testing laboratory.
-[^1] A visual of this implementation, showing a screenshot of the digital twin of the METU high voltage testing laboratory on the left, and a photograph of the actual site on the right, can be accessed at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660.
+Furthermore, in recent years, digital twins have also gained importance in the monitoring of processes in industrial environments, especially in hazardous or hard-to-reach environments or when it is impractical to be on site for monitoring.
+[^1] A screenshot of the digital twin of the METU high voltage testing laboratory is available at https://cng-eee.metu.edu.tr/wp-content/Digital_twin.pdf?_t=1726991660.
 The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
-Such systems, involving real-time monitoring, training, etc., often require computation that relies on data sent from remote locations to maintain good performance.
-Motivated by this, we focus on goal-oriented communication design for the remote inference problem [4, 5].
-Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks [6].
-This concept emphasizes that networks should focus on the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source.
+We focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [5, 6], which are often essential for such applications.
+Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks.
+This concept emphasizes that networks can facilitate the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source.
 Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8].
@@
 Recent efforts in the communication and control communities highlight the value of this approach in reducing communication requirements while maintaining application performance [9, 10, 11, 12, 13, 14].
-Since we cannot expect the link layer to operate jointly with the application layer for all possible applications, a surrogate metric that the link layer can utilize when deciding which samples to select and transmit is useful.
-The Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet.
-We utilize the AoI as an intermediate metric that can be monitored by the link layer, independent of the specific application.
-A mapping of application performance in relation to the AoI is all that is required for the link layer to function in a goal-oriented manner.
-This work builds upon the findings presented in the series of papers [4, 16, 5, 17, 1], which first demonstrated the effectiveness of the AoI metric as a surrogate for enhancing communication in the remote tracking or inference of a process.
-In this paper, we study a remote inference problem, where an intelligent model infers the real-time value of a target signal based on data transmitted from a remote location.
-We assume that the transmitter regularly samples a source and forms data packets containing a number of data samples, which we refer to as the packet length, before each transmission.
-The intelligent model on the receiver side performs inference using these data packets.
-The inference error is a function of the AoI and the length of the data packet used by the intelligent model.
-Recent studies [4, 5] have shown that, in various remote inference problems, the inference error (4) does not always increase monotonically with AoI.
-This contradicts the conventional assumption that a packet with AoI $=0$ necessarily yields better performance than one with AoI $>0$.
-Therefore, rather than the generate-at-will model [18, 19], we adopt the “selection-from-buffer” model proposed in [4, 5], which enables the scheduler to choose either fresh or stale data samples from the buffer.
-Moreover, the inference error decreases monotonically as the packet length increases [17].
-This improvement comes at the cost of higher resource usage and increased delay.
-In our communication design, we explore both the possibly non-monotonic relationship between inference error and AoI, as well as the interplay between packet length and transmission delay.
-Due to the need for extreme connectivity, next-generation communication networks are expected to involve numerous connections, leading to the existence of multiple routes between any nodes in the network.
-The relays in such networks may enforce specific routes for certain data flows, which can dynamically change over time to meet the varying demands of multiple flows.
+Age of Information (AoI) at time $t$, denoted by $\Delta(t)$, is defined in [15] as $\Delta(t)=t-U_{t}$, where $U_{t}$ is the generation time of the most recently delivered data packet.
+The application-layer real-time estimation and inference performance is mapped as a function of the AoI.
+This function is all that is required for the link layer to operate in a goal-oriented manner.
+This work builds upon the findings presented in a series of papers [5, 16, 6, 17, 1, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
+In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$, based on a data packet, $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$, transmitted from a remote location.
+The data packet $X_{t-\delta}^{l}$ comprises samples of the source signal $V_{t}$ at the transmitter.
+Here, $\delta$ represents the AoI of the freshest sample in the packet, and $l$ denotes the packet length, corresponding to the number of samples included in the packet.
+The inference error is characterized as a function of both the AoI and the packet length.
+Recent studies [5, 6] have demonstrated that, for various remote inference systems, the relationship between the inference error and AoI for a given packet length can be non-monotonic.
+This finding contradicts the conventional assumption that a packet with ${\text{AoI}=0}$ necessarily yields better performance than one with ${\text{AoI}>0}$.
+Consequently, rather than adopting the generate-at-will model [21, 18], we use the “selection-from-buffer” model proposed in [5, 6], which enables the scheduler to choose either fresh or stale data samples from the buffer.
+Such non-monotonicity between the inference error and AoI arises when the source signal, $V_{t}$, and the target signal, $Y_{t}$, exhibit a delayed relationship, such as $Y_{t}=f(V_{t-\Upsilon})$ [6, 22, 20].
+For instance, $\Upsilon$ may represent a combination of the communication delay between the controller and the actuator and the actuation delay in a physical networked control system.
+An example of this scenario is illustrated in the leader-follower robot setup in [6, Fig.
+2].
+Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [6, Fig.
+4].
+In such cases, when communication delays prevent the transmission of sufficiently fresh samples to the receiver, transmitting aged samples can exploit the periodicity to enhance performance.
+Moreover, the inference error is non-increasing with packet length for a given AoI [17], as additional information can only improve the estimation accuracy.
+However, this improvement comes at the cost of longer transmission delays.
+In our communication design, we investigate the interplay between packet length and transmission delay, as well as the possibly non-monotonic relationship between inference error and AoI.
+Finally, next-generation communication networks are expected to accommodate numerous connections, resulting in multiple routes between any pair of nodes.
+The relays in such networks may enforce specific routes for certain data flows, which can dynamically change over time to address the varying demands of multiple flows.
 For instance, consider a network involving both terrestrial and non-terrestrial connections.
-A flow can be fully serviced by non-terrestrial connections, terrestrial connections, or a combination of both, depending on conditions such as the priority of the flow, network congestion, or the availability of satellites.
-Another example is busy industrial environments, where communication routes may be intermittently unavailable due to shadowing from moving objects, causing flows to be rerouted over time.
-Consequently, goal-oriented communication design for remote inference must be adaptable to delay conditions significantly varying with memory.
+A flow may be serviced entirely by non-terrestrial connections, terrestrial connections, or a combination of both, depending on factors such as the flow’s priority, network congestion, or satellite availability.
+Therefore, goal-oriented communication design must be adaptable to delay conditions that vary significantly with memory.
 To that end, the technical contributions of this paper are:
-  • Extending the system models in [5, 17], we formulate and solve a learning and communication co-design problem for remote inference under two-way delay that varies significantly with memory.
-We consider both time-invariant and time-variable packet length selection scenarios to address the varying computational capabilities of practical systems.
-The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference system, thereby making the communication goal-oriented.
+  • We extend the system models in [6, 17] to more practical scenarios by considering Markovian delay on both forward and feedback links.
+Based on this extension, we formulate and solve a learning and communication co-design problem for real-time remote inference.
+We consider both time-invariant and time-variable packet length selection to address the varying computational capabilities of practical systems.
+The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
   • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP).
 Such problems are often solved using dynamic programming [23, 24] and typically do not have closed-form solutions.
-However, we derive a closed-form solution for this scenario, presented in Theorem 1.
+However, we derive a closed-form solution to this problem for any constant packet length, presented in Theorem 1.
+The solution consists of two main parts: (i) the AoI of data packets to be sent are determined as a stationary function of the network state, i.e., the delay state, and (ii) the transmission instants are governed by an index-based threshold policy.
+The index function is expressed in terms of the delay state of the network and the AoI at the receiver.
+We then separately optimize the value of constant packet length.
   • We also model the learning and communication co-design problem under time-variable packet length selection as an infinite-horizon average-cost SMDP.
@@
 Solving the simplified Bellman equation using dynamic programming has significantly lower time complexity compared to solving the original Bellman optimality equation.
-  • We conduct two experiments to evaluate the performance of the derived optimal scheduling policies: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which offers a trace-driven evaluation.
-Simulation results indicate the substantial benefits of adopting a goal-oriented communication strategy for remote inference.
+  • We conduct two experiments to evaluate the performance of our optimal scheduling policies: (i) remote inference of auto-regressive (AR) processes, offering a model-based evaluation; and (ii) cart-pole state prediction, providing a trace-driven evaluation.
+In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
 I-A Related Work
@@
 The works most closely related to this paper are [6, 17], which developed scheduling policies for remote inference, considering a possibly non-monotonic dependency between AoI and practical performance.
-While [5] developed scheduling policies for fixed packet length, the paper [17] jointly optimized packet length and scheduling strategies.
-Both the paper [5, 17] considered random IID delay for packet transmissions from the transmitter to the receiver and assumed delay-free feedback from the receiver to the transmitter.
-In practical systems, a non-zero feedback delay exists [40, 41], which affects the performance of the remote inference algorithm.
-Moreover, due to the envisioned structure of next-generation communication networks, which contain multiple routes between any pair of nodes, the IID delay may not be a good model.
-This necessitates a model that captures the memory in the delay process, as in [19].
-Extending the problem formulations in [5, 17], this paper designs a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay significantly varying with memory.
+While [6] developed scheduling policies for fixed packet length, [17] jointly optimized packet length and scheduling strategies.
+However, both studies [6, 17] assumed random i.i.d.
+delay for packet transmissions and immediate feedback.
+This paper extends the problem formulations in [6, 17] to more practical scenarios and develops a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay (i.e., incorporating random feedback delay [40, 41]) that varies significantly with memory.
 II System Model
-We consider a remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig.
+Figure 1: A remote inference system that adopts the selection-from-buffer medium access model [6, 5].
+We consider a time-slotted remote inference system, which comprises a data source, a transmitter, and a receiver, as illustrated in Fig.
 1.
-On the transmitter side, the source signal $V_{t}$ is regularly sampled from the data source, and the samples are stored in a buffer containing the most recent $B$ samples.
-The scheduler determines when to transmit and which samples in the buffer to transmit.
-On the receiver side, a predictor, e.g., a pre-trained neural network, infers the real-time value of a target signal $Y_{t}$ based on the most recently received samples.
-Such remote inference systems are crucial for many real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and Cyber-Physical Systems (CPS), where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
-Figure 1: A remote inference system, where the packet $X_{t-b}^{l}$ represents the group of samples $\big{(}V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1}\big{)}$.
-We assume that the system is time-slotted, and it starts operation at time slot $t=0$.
-We adopt the selection-from-buffer model proposed in [5], where, at each time slot $t$, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer while the oldest sample is discarded.
+On the transmitter side, the source signal $V_{t}$ is regularly sampled in each time slot $t$.
+The samples are stored in a buffer containing the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$, waiting to be sent to the receiver.
+On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of a target signal $Y_{t}$ of interest.
+In practical time-series forecasting algorithms, the predictor needs to use multiple consecutive samples, e.g., $X_{t-b}^{l}=(V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1})$, to infer $Y_{t}$.
+We will design a transmission scheduler that determines (i) when to send a group of consecutive samples over the communication network to the receiver and (ii) which group of consecutive samples in the buffer to send.
+This medium access model is called “selection-from-buffer,” which was proposed recently in [5, 6].
+Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
+We assume that the system starts operation at time slot $t=0$.
+In the selection-from-buffer model, a new sample of the source signal $V_{t}\in\mathcal{V}$ is added to the buffer at the beginning of each time slot $t$, and the oldest sample is discarded.
 The buffer, with size $B$, is assumed to be initially full, storing the samples $\big{(}V_{0},V_{-1},\ldots,V_{-B+1}\big{)}$.
-Thus, at any time slot $t$, the buffer contains the most recent $B$ samples, $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$.
-Before the $i$-th transmission, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{t-b_{i}}^{l_{i}}=\big{(}V_{t-b_{i}},V_{t-b_{i}-1},\ldots,V_{t-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l}$.
-Here, $l_{i}\in\{1,\ldots,B\}$ and $b_{i}\in\{0,\ldots,B-l_{i}\}$ represents the packet length and buffer position of the $i$-th packet $X_{t-b_{i}}^{l_{i}}$, respectively.
-Fig.
-2 indicates a sample evolution for two consecutive transmission periods.
-The packet $X_{t-b_{i}}^{l_{i}}$ is submitted at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$.
+Thus, at any time slot $t$, the buffer contains the most recent $B$ samples $\big{(}V_{t},V_{t-1},\ldots,V_{t-B+1}\big{)}$.
+The transmitter sends status update packets to the receiver one by one.
+Let $S_{i}$ and $D_{i}$ represent the transmission start time and delivery time of the $i$-th transmitted packet, respectively, such that $S_{i}<D_{i}$.
+At time slot $t=S_{i}$, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{S_{i}-b_{i}}^{l_{i}}=\big{(}V_{S_{i}-b_{i}},V_{S_{i}-b_{i}-1},\ldots,V_{S_{i}-b_{i}-l_{i}+1}\big{)}\in\mathcal{V}^{l_{i}}$.
+Here, $l_{i}\in\{1,\ldots,B\}$ represents the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$ while $b_{i}$ represents the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer.
+When $B=1$, the selection-from-buffer model reduces to the generative-at-will model which has been studied extensively in previous works [37, 18, 38, 39, 19, 27, 28, 29].
+Hence, the packet $X_{S_{i}-b_{i}}^{l_{i}}$ is submitted to the communication network at time slot $S_{i}$ and delivered to the receiver at time slot $D_{i}$.
 Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$.
 We assume that the transmitter always waits for feedback before submitting a new packet.
-In other words, the transmitter remains silent during the time slots between $S_{i}$ and $A_{i}$ for all $i$.
+In other words, the transmitter remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$.
+Fig.
+2 provides an illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
 II-A Communication Network Model
-The remote inference system operates on top of a communication network with multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver.
-Each packet submitted into the network is directed to a route based on the network status regarding availability and congestion.
+Figure 2: An illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
+Consider that the communication network may have multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver.
+The network directs each submitted packet to one of these available routes.
 If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route.
+Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$.
+Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay.
+Additionally, let $C_{i}$ represent the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization.
+The delay state $C_{i}$ is determined by the route used in the $i$-th packet transmission epoch.
 Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length.
 The delay is at least one time slot for each packet or ACK transmission.
-Let the $i$-th epoch consist of the time slots in the interval $[A_{i-1},A_{i})$.
-Define $C_{i}$ as the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization.
-The delay state $C_{i}$ represents the network route used in the $i$-th epoch and specifies the transmission and feedback delay distributions.
 We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$.
 The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise.
-Let $T_{i}(l_{i})=D_{i}-S_{i}$ and $F_{i}=A_{i}-D_{i}$ represent the transmission and feedback delays incurred in the $i$-th epoch, respectively.
-The receiver detects the delay state $c_{i}$ at each delivery time $D_{i}$ and embeds this information into the ACK message.
-The ACK notifies the scheduler that the previous packet transmission is complete and that the next packet can be submitted to the network.
-The network is reliable, meaning no packet or message is lost during transmission.
-II-B Inference Error
-The Age of Information (AoI) on the receiver side, $\Delta(t)$, is the time difference between the current time $t$ and the generation time $t-\Delta(t)$ of the freshest sample $V_{t-\Delta(t)}$ in the most recently delivered packet $X_{t-\Delta(t)}^{l}$.
+The receiver detects the delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message.
+The ACK notifies the transmitter of (i) the delay state experienced during the previous packet transmission, (ii) the completion of the previous packet transmission, and (iii) that the next packet can be submitted to the network.
+The network is reliable, meaning no packet is lost during transmission.
+II-B Inference Error as a Function of AoI and Packet Length
+Age of Information (AoI) on the receiver side, $\Delta(t)$, is the time difference between the current time $t$ and the generation time $t-\Delta(t)$ of the freshest sample $V_{t-\Delta(t)}$ in the most recently delivered packet $X_{t-\Delta(t)}^{l}$ [15].
 The AoI at time slot $t$ is determined by
 $\Delta(t)=t-S_{i}+b_{i},\text{ if }D_{i}\leq t<D_{i+1}.$ (1)
-Figure 2: A sample evolution for two consecutive transmission periods.
-Let $d(t)$ represent the length of the most recently delivered packet by time slot $t$, which can be written as follows:
+Let $d(t)$ represent the sequence length of the most recently delivered packet by time slot $t$, which is given by
 $d(t)=l_{i},\text{ if }D_{i}\leq t<D_{i+1}.$ (2)
-We focus on a class of widely used supervised learning algorithms known as Empirical Risk Minimization [42].
-We assume that, for each packet length $l\in\{1,2,\ldots,B\}$, the receiver contains a trained neural network $\phi^{*}_{l}\big{(}\cdot\big{)}$ that takes the AoI $\Delta(t)\in\mathbb{Z}^{+}$ and the packet $X_{t-\Delta(t)}^{l}\in\mathcal{V}^{l}$ as inputs and produces an output $a\in\mathcal{A}$ to infer the real-time value of the target signal $Y_{t}\in\mathcal{Y}$.
-The performance is evaluated through a loss function $L\big{(}\cdot\big{)}:\mathcal{Y}\times\mathcal{A}\mapsto\mathbb{R}$.
-The incurred loss is $L(y,a)$ when the output $a$ is taken to predict $Y_{t}=y$.
-Given the AoI $\delta$ and packet length $l$, the training problem to obtain the neural network $\phi^{*}_{l}\big{(}\cdot,\cdot\big{)}$ is formulated as:
-$\min_{\phi\big{(}\cdot,\cdot\big{)}\in\Lambda}\mathbb{E}_{Y,X^{l}\sim P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}}\big{[}L\big{(}Y,\phi\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (3)
-where $\Lambda$ is the set of all mappings that the neural network can generate, and $P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}$ represents the empirical distribution of the target signal $\tilde{Y}_{0}$ and the packet $\tilde{X}_{-\delta}^{l}$.
-The AoI $\delta$ is the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}$.
-These trained predictors are used to estimate the real-time value of the target signal $Y_{t}$ on the receiver side.
+We assume that the predictor employs a supervised learning algorithm based on Empirical Risk Minimization [42].
+For each packet length $l\in\{1,2,\ldots,B\}$, the receiver contains a pre-trained neural network $\phi^{*}_{l}:\mathbb{Z}^{+}\times\mathcal{V}^{l}\mapsto\mathcal{A}$, which takes the AoI $\Delta(t)$ and the packet $X_{t-\Delta(t)}^{l}$ as inputs and produces an action $a$ to infer the real-time value of the target signal $Y_{t}$.
+The performance is evaluated through a loss function ${L:\mathcal{Y}\times\mathcal{A}\mapsto\mathbb{R}}$.
+The incurred loss is $L(y,a)$ when the action $a$ is taken to predict $Y_{t}=y$.
+Given the AoI $\delta$ and packet length $l$, the training problem to obtain the neural network $\phi^{*}_{l}$ is formulated as:
+$\phi^{*}_{l}=\operatorname*{arg\,min}_{\phi\in\Lambda}\mathbb{E}_{Y,X^{l}\sim P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}}\big{[}L\big{(}Y,\phi\big{(}\delta,X^{l}\big{)}\big{)}\big{]},$ (3)
+where $\Lambda$ is the set of all mappings that the neural network can generate, and $P_{\tilde{Y}_{0},\tilde{X}_{-\delta}^{l}}$ represents the empirical distribution of the target signal $\tilde{Y}_{0}$ and the packet $\tilde{X}_{-\delta}^{l}$ in the training dataset.
+The AoI $\delta$ denotes the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}^{l}$.
+These trained neural networks are used to predict the real-time value of the target signal $Y_{t}$ on the receiver side.
 We assume that the process ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ is stationary, and that the processes ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ and ${\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}}$ are independent.
@@
 where $P_{Y_{t},X_{t-\delta}^{l}}$ represents the joint distribution of the target signal $Y_{t}$ and the packet $X_{t-\delta}^{l}$.
-We are interested in loss functions $L\big{(}\cdot\big{)}$ for which the inference error (4) is bounded.
-Apart from this restriction, the loss function $L\big{(}\cdot\big{)}$ can be chosen based on the goal of the remote inference system.
+The loss function $L$ can be chosen based on the goal of the remote inference application.
 For example, a quadratic loss function $L_{2}\big{(}y,\hat{y}\big{)}=\lVert y-\hat{y}\rVert_{2}^{2}$ is used in neural network-based minimum mean-squared estimation, where the action $a=\hat{y}$ is an estimate of the target signal $Y_{t}=y$ and $\lVert y\rVert_{2}$ is the Euclidean norm.
@@
 For a given AoI $\delta$, the inference error is a non-increasing function of packet length [17].
-However, for a given packet length $l$, the inference error does not always increase monotonically with the AoI [4, 5].
-This observation contradicts the conventional assumption that fresher packets yield smaller inference errors.
-Theoretical and experimental explanations of this counter-intuitive observation was provided in [4, 5].
-Therefore, we do not adopt the generate-at-will model [18, 19] in this paper.
-II-C Structure of a Scheduling Policy
-Upon the ACK reception time $A_{i}$, the scheduler determines a waiting time $\tau_{i+1}$ to specify the next submission time slot $S_{i+1}=A_{i}+\tau_{i+1}$.
-Then, a data packet $X_{S_{i+1}-b_{i+1}}^{l_{i+1}}$ is formed by selecting the packet length $l_{i+1}$ and buffer position $b_{i+1}$.
-A scheduling policy $\pi\in\Pi$ is defined as a tuple $\pi=\big{(}f,\ell,g\big{)}$, where $f=\big{(}b_{2},b_{3},\ldots\big{)}$ is the buffer position sequence, $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$ is the packet length sequence, $g=\big{(}\tau_{2},\tau_{3},\ldots\big{)}$ is the waiting time sequence, and $\Pi$ denotes the set of all causal policies $\pi$.
-We assume that the scheduling policy does not employ the knowledge of the observed process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$.
-This assumption implies that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are independent.
+On the other hand, for a given packet length $l$, the inference error does not always increase monotonically with the AoI [5, 6].
+When minimizing non-decreasing functions of AoI, fresher data packets are preferable, and the generate-at-will medium access model works well.
+However, to minimize inference error, which is not necessarily non-decreasing with respect to AoI, the selection-from-buffer model—allowing for the choice between fresh and older data packets—is more effective.
+II-C Structure of the Scheduling Policy
+Upon receiving the ACK at time slot $A_{i}$, the scheduler determines a waiting time $\tau_{i+1}$ to specify the next submission time slot as $S_{i+1}=A_{i}+\tau_{i+1}$.
+Then, a new data packet $X_{S_{i+1}-b_{i+1}}^{l_{i+1}}$ is formed with packet length $l_{i+1}$ and relative buffer position $b_{i+1}$.
+A scheduling policy $\pi\in\Pi$ is defined as a tuple $\pi=\big{(}f,\ell,g\big{)}$, where $f=\big{(}b_{2},b_{3},\ldots\big{)}$ is the relative buffer position sequence, $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$ is the packet length sequence, $g=\big{(}\tau_{2},\tau_{3},\ldots\big{)}$ is the waiting time sequence.
+In addition, let $\Pi$ denote the set of all causal policies $\pi$.
+We assume that the scheduler does not utilize knowledge of the process $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$.
+This assumption implies that the processes $\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}$ and $\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}$ are mutually independent.
 The initial conditions of the system are assumed to be $S_{1}=0$, $l_{1}=1$, $b_{1}=0$, $c_{1}\in\{1,2,\ldots,N\}$, and $\Delta(0)$ is a finite constant.
@@
 While designing goal-oriented communication protocols for remote inference, selecting a constant packet length is a simple and often effective choice.
-If packet lengths vary dynamically from one packet to the next, the receiver needs to reconfigure the estimator over time to accommodate the changing input dimensions, which adds complexity.
-In this section, we consider systems where this complexity is avoided by choosing a time-invariant packet length sequence.
-Systems that can vary the packet length over time are discussed in Section IV.
+If packet lengths vary dynamically from one packet to the next, the receiver needs to reconfigure the predictor over time to accommodate the changing input dimensions, which adds complexity.
+This section considers systems where this complexity is avoided by choosing a time-invariant packet length sequence.
+Systems that can vary the packet length over time will be discussed later in Section IV.
 III-A Co-design Problem Formulation
@@
 $\Pi_{l}=\{\pi\in\Pi:l=l_{2}=l_{3}=\ldots\}.$ (5)
-The learning and communications co-design problem with time-invariant packet length selection can be formulated as a two-layer nested optimization problem:
+The packet length affects both learning and communication.
+If it is selected based solely on learning performance, communication may suffer since transmission delays increase with packet length.
+Therefore, we solve a learning and communications co-design problem under time-invariant packet length selection, formulated as the following two-layer nested optimization problem.
 $\varepsilon_{l,\text{opt}}=\inf_{\pi\in\Pi_{l}}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{T-1}\varepsilon\big{(}\Delta(t),l\big{)}\Bigg{]},$ (6)
@@
 where $\varepsilon_{l,\text{opt}}$ is the optimum value of the inner optimization problem (6) for a given packet length $l$, and $\varepsilon_{\text{opt}}^{\text{inv.}}$ is the optimum value of the two-layer nested optimization problem (6)-(7) with the optimal time-invariant packet length.
-The inner optimization problem (6) can be cast as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP).
-While dynamic programming algorithms, namely policy iteration and value iteration, are typically employed to solve such problems [20, 21], we are able derive a closed-form solution that achieves significantly lower computational complexity than dynamic programming.
-Moreover, the outer optimization problem (7) is a straightforward search over the integer values $\{1,2,\ldots,B\}$.
-Unlike the formulation in [17], this problem formulation accounts for varying delay conditions between the transmitter and the receiver and does not assume immediate feedback.
+The inner problem (6) optimizes the relative buffer position sequence $f$ and waiting time sequence $g$ for a given packet length $l\in\{1,2,\ldots,B\}$.
+This problem can be cast as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP).
+Although dynamic programming algorithms, such as policy iteration and value iteration, are typically used to solve such problems [23, 24], we are able to derive a closed-form solution that achieves significantly lower computational complexity than dynamic programming.
+Moreover, the outer problem (7) optimizes the packet length and involves a straightforward search over the integer values $\{1,2,\ldots,B\}$.
+Problem (6)-(7) is a more general formulation than that considered in [17, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in multi-route communication networks.
 III-B Optimal Solution to (6)-(7)
-We define each ACK reception time $A_{i}$ as a decision time for the SMDP (6).
-The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch.
-The actions are the waiting time $\tau_{i+1}$ and the buffer position $b_{i+1}$.
-A detailed description of the SMDP can be found in Appendix A.
-The Bellman optimality equation of this SMDP can be formulated as:
+Each time slot $A_{i}$ is a decision time for the SMDP (6).
+The system state at each $A_{i}$ is given by a tuple $\big{(}\Delta(A_{i}),C_{i}\big{)}$, where $\Delta(A_{i})$ represents the AoI at time-slot $A_{i}$, and $C_{i}$ indicates the delay state in the $i$-th packet transmission epoch.
+A realization of the system state is denoted by $\big{(}\delta,c_{i}\big{)}$.
+The actions determined at time slot $A_{i}$ are the waiting time $\tau_{i+1}$ and the buffer position $b_{i+1}$.
+A detailed description of this SMDP can be found in Appendix A.
+The Bellman optimality equation of the SMDP (6) is formulated as
 $h_{1}\big{(}\delta,c_{i}\big{)}=\min_{\begin{subarray}{c}\tau_{i+1}\in\{0,1,\ldots\}\\ b_{i+1}\in\{0,1,\ldots,B-l\}\end{subarray}}$
@@
 $+\mathbb{E}\big{[}h_{1}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]},$ (8)
-where $h_{1}(\delta,c_{i})$ is the relative value function.
+where $\delta\in\mathbb{Z}^{+}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{1}(\delta,c_{i})$ is the relative value function.
 We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6).
@@
 $\gamma\big{(}\delta,d,l,c\big{)}=\min_{\tau\in\{1,2,\ldots\}}\frac{1}{\tau}\sum_{k=0}^{\tau-1}\mathbb{E}\bigg{[}\varepsilon\big{(}\delta+T_{i+1}(l)+k,d\big{)}\bigg{|}C_{i}=c\bigg{]},$ (9)
-for all $\delta\in\{1,2,\ldots\}$, $d,l\in\{1,2,\cdot,B\}$, and $c\in\{1,2,\ldots,N\}$.
+for all $\delta\in\mathbb{Z}^{+}$, $d,l\in\{1,2,\ldots,B\}$, and $c\in\{1,2,\ldots,N\}$.
 Theorem 1.
@@
 $\mathfrak{D}=F_{i+1}+\tau_{l}\big{(}b+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, and $\gamma\big{(}\cdot\big{)}$ is defined in (9). Furthermore, $\beta$ is the unique root of
-$\mathbb{E}\Bigg{[}\sum_{t=A_{i}(\beta)}^{A_{i+1}(\beta)-1}\varepsilon\big{(}\Delta(t)\big{)}\Bigg{]}-\beta\mathbb{E}\big{[}A_{i+1}(\beta)-A_{i}(\beta)\big{]}=0,$ (12)
-and it is exactly the optimum value $\varepsilon_{l,\text{opt}}$ of problem (6).
+$\mathbb{E}\Bigg{[}\sum_{t=A_{i}(\beta)}^{A_{i+1}(\beta)-1}\varepsilon\big{(}\Delta(t),l\big{)}\Bigg{]}-\beta\mathbb{E}\big{[}A_{i+1}(\beta)-A_{i}(\beta)\big{]}=0,$ (12)
+and $\beta=\varepsilon_{l,\text{opt}}$ is exactly the optimum value of problem (6).
+$A_{i}(\beta)$ in equation (12) represents the $i$-th ACK reception time slot when the scheduling policy is determined according to the rules specified by equations (10) and (1) for a given $\beta$.
 Proof Sketch.
 We prove Theorem 1 in three steps:
-Step 1: We obtain two key results by using the Bellman optimality equation (III-B):
-  • The optimal waiting time $\tau_{i+1}^{*}$ is determined by the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ given by equation (10) for all $i$.
+Step 1: We first establish two key results using the Bellman optimality equation (III-B):
+  • The optimal waiting time $\tau_{i+1}^{*}$ is determined by the index-based threshold rule
+    $\min\{k\geq 0:\gamma\big{(}\delta+k,l,l,c_{i}\big{)}\geq\varepsilon_{l,\text{opt}}\}$
+    for all $i$. Here, $\gamma\big{(}\cdot\big{)}$ is the index function defined in (9).
   • The delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining the optimal buffer position $b_{i+1}^{*}$ for all $i$.
@@
 $+\mathbb{E}\bigg{[}h_{1}^{\prime}\big{(}C_{i+1}\big{)}\bigg{|}C_{i}=c_{i}\bigg{]},$ (13)
-where $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$, and $h_{1}^{\prime}\big{(}c_{i}\big{)}$ is the relative value function.
-The Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision-epoch optimization problem, allowing any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ to be expressed as shown in (1).
-Step 3: We determine that the optimal value $\varepsilon_{l,\text{opt}}$ for the inner optimization problem (6) is the unique root of (12).
+where $c_{i}\in\{1,2,\ldots,N\}$, $h_{1}^{\prime}\big{(}c_{i}\big{)}$ is the relative value function, and $\vartheta_{i+1}=F_{i+1}+\tau_{l}\big{(}b_{i+1}+T_{i+1}(l)+F_{i+1},C_{i+1}\big{)}+T_{i+2}(l)$.
+Only the first term on the right-hand side of equation (III-B) depends on the action $b_{i+1}$.
+Thus, the Bellman optimality equation (III-B) is decomposable and can be solved as a per-decision-epoch optimization problem, allowing any $b_{i+1}^{*}$ in the optimal buffer position sequence ${f^{*}=\big{(}b_{2}^{*},b_{3}^{*},\ldots\big{)}}$ to be expressed as shown in equation (1).
+Step 3: Finally, we show that the optimal value $\varepsilon_{l,\text{opt}}$ of problem (6) is the unique root of equation (12).
 The detailed proof is provided in Appendix B. ∎
-The optimal scheduling policy for the inner optimization problem (6), outlined in Theorem 1, is well-structured.
+The optimal scheduling policy for the inner optimization problem (6), as outlined in Theorem 1, is well-structured.
 Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$.
-The index function $\gamma\big{(}\cdot\big{)}$, provided in (9), can be readily calculated for any state $\big{(}\delta,c_{i}\big{)}$, using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$.
-Furthermore, according to equation (1), $b_{i+1}^{*}$ is designed to minimize the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10).
-Also, each optimal buffer position $b_{i+1}^{*}$ depends solely on the delay state $c_{i}$ in the $i$-th epoch, and is independent of $\delta=\Delta(A_{i})$.
-This result can be better understood by deeply analyzing the AoI evolution.
-At each delivery time slot $D_{i+1}$, the AoI resets to a realization of $b_{i+1}^{*}+T_{i+1}(l)$.
-Due to these regular resets, the AoI evolution between consecutive delivery slots $D_{i+1}$ and $D_{i+2}$ is independent of the AoI evolution outside this interval.
-Moreover, $b_{i+1}^{*}$ is a parameter to control the AoI evolution between $D_{i+1}$ and $D_{i+2}$.
+The index function $\gamma\big{(}\cdot\big{)}$, defined in equation (9), can be readily computed for any state $\big{(}\delta,c_{i}\big{)}$ using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$.
+Although the index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ closely resembles the waiting time rule provided in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$.
+Because problem (6)-(7) involves non-zero feedback delay and time-varying delay distribution, the existence of such a simple waiting time rule was not evident and required additional technical efforts for its derivation.
+Furthermore, according to equation (1), the optimal buffer position $b_{i+1}^{*}$ is designed by minimizing the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10).
+Each $b_{i+1}^{*}$ depends solely on the delay state $C_{i}=c_{i}$ in the $i$-th epoch and is independent of the AoI state $\delta=\Delta(A_{i})$.
+For the case of i.i.d.
+transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [17, Theorem 1].
+However, when the feedback delay is non-zero and the delay state $C_{i}$ changes according to a Markov chain, achieving optimal performance requires that the optimal buffer position $b_{i+1}^{*}$ is a function of the delay state $C_{i}=c_{i}$, and this function is exactly given by (1).
 Finally, the optimum value $\varepsilon_{l,\text{opt}}$, which is the threshold in the waiting time rule (10) and necessary for solving (1), is the unique root of equation (12).
-This equation can be efficiently solved using low-complexity algorithms [25, Algorithms 1-3], such as bisection search.
-The optimal policy in Theorem 1 has two key distinctions compared to the policy established in [17, Theorem 1] for remote inference under time-invariant packet length selection, IID transmission delays, and immediate feedback.
-First, although the threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ in (10) closely resembles the waiting time rule derived in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state.
-For the problem formulation (6)-(7), which involves a finite number of delay states with non-zero feedback delay, the existence of such a waiting time rule was not evident and required additional technical efforts for its derivation.
-Second, in [17, Theorem 1], the optimal buffer position sequence is time-invariant due to the assumption of IID transmission delays.
-However, when the delay state changes with memory, achieving optimal performance requires a buffer position sequence that can adapt to variations in delay statistics.
-Theorem 1 unveils that the optimal buffer position sequence is determined by solving $N$ per-sample optimization problems, each related to a distinct delay state, without requiring pre-computation of the relative value function $h_{1}\big{(}\cdot\big{)}$.
-We can solve the inner optimization problem (6) for any given packet length $l$ by using Theorem 1.
-Then, the outer optimization problem (7) is just a straightforward search over the integer values $\{1,2,\ldots,B\}$, and $\varepsilon_{\text{opt}}^{\text{inv.}}=\varepsilon_{l^{*},\text{opt}}$.
+This equation can be efficiently solved using low-complexity algorithms [26, Algorithms 1-3], such as bisection search.
+After solving the inner optimization problem (6) using Theorem 1, the outer optimization problem (7) is just a straightforward search over the integer values $\{1,2,\ldots,B\}$.
 IV Learning and Communications Co-design: Time-Variable Packet Length Selection
-In contrast to the previous case, where high computational complexity is avoided, here we aim to achieve optimal performance without any restrictions on the packet length sequence $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$.
+In contrast to the previous case of time-invariant packet length, this section aims to achieve optimal performance without imposing restrictions on the packet length sequence $\ell=\big{(}l_{2},l_{3},\ldots\big{)}$.
 This unconstrained approach allows us to maximize the system’s potential by dynamically adjusting the packet length $l_{i}$ over time.
@@
 IV-A Co-design Problem Formulation
-The learning and communications co-design problem with time-variable packet length selection is formulated as:
+The learning and communications co-design problem with general, time-variable packet length selection is formulated as
 $\varepsilon_{\text{opt}}=\inf_{\pi\in\Pi}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi}\Bigg{[}\sum_{t=0}^{T-1}\varepsilon\big{(}\Delta(t),d(t)\big{)}\Bigg{]},$ (14)
@@
 Problem (14) is an infinite-horizon average-cost SMDP, which is more challenging to solve compared to problem (6).
-We formulate a Bellman optimality equation for this SMDP and then simplify it using a structural result regarding the optimal policy.
+We will first formulate a Bellman optimality equation for this SMDP and then simplify it using a structural result regarding the optimal policy.
 The time complexity for solving the simplified Bellman optimality equation using dynamic programming algorithms is substantially lower than that for solving the original Bellman optimality equation.
 IV-B Optimal Solution to (14)
-Each ACK reception time $A_{i}$ is a decision time for the SMDP (14).
-The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,d,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch.
-The actions to be taken at each decision time $A_{i}$ are the waiting time $\tau_{i+1}$, the packet length $l_{i+1}$, and the buffer position $b_{i+1}$.
-A detailed description of the SMDP can be found in Appendix C.
-The Bellman optimality equation of this SMDP can be expressed as:
+In the SMDP (14), the system state at each decision time $A_{i}$ is given by the tuple $\big{(}\Delta(A_{i}),d(A_{i}),C_{i}\big{)}$, where $\Delta(A_{i})$ denotes the AoI at time slot $A_{i}$, $d(A_{i})$ indicates the length of the most recently received packet by that time, and $C_{i}$ represents the delay state in the $i$-th packet transmission epoch.
+A realization of the system state is denoted by $\big{(}\delta,d,c_{i}\big{)}$.
+The actions taken at each decision time $A_{i}$ include the waiting time $\tau_{i+1}$, the packet length $l_{i+1}$, and the buffer position $b_{i+1}$.
+A detailed description of this SMDP can be found in Appendix C.
+The Bellman optimality equation for the SMDP (14) can be expressed as
 $h_{2}\big{(}\delta,d,c_{i}\big{)}=\min_{\begin{subarray}{c}\tau_{i+1}\in\{0,1,\ldots\}\\ l_{i+1}\in\{1,2,\ldots,B\}\\ b_{i+1}\in\{0,1,\ldots,B-l_{i+1}\}\end{subarray}}$
@@
 $+\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]},$ (15)
-where $h_{2}\big{(}\delta,d,c_{i}\big{)}$ is the relative value function.
-By solving the Bellman optimality equation (IV-B) using dynamic programming, we can obtain a solution to problem (14).
+where $\delta\in\mathbb{Z}^{+}$, $d\in\{1,2,\ldots,B\}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{2}\big{(}\delta,d,c_{i}\big{)}$ is the relative value function.
+By solving the Bellman optimality equation (IV-B) using dynamic programming, we can obtain a solution to (14).
 The time complexity of this dynamic programming algorithm is high, as it requires the joint optimization of three variables for each state $\big{(}\delta,d,c_{i}\big{)}$.
-Theorem 2 provides a threshold rule for determining the waiting time action at each decision time $A_{i}$ and simplifies the Bellman optimality equation.
-The time complexity for solving the simplified Bellman optimality equation is significantly lower than that for solving (IV-B).
+The following Theorem 2 provides a threshold rule for determining the waiting time action at each $A_{i}$, simplifying the Bellman optimality equation.
 Theorem 2.
@@
 where $\gamma\big{(}\cdot\big{)}$ is defined in (9).
-Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the simplified Bellman optimality equation
+Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the following simplified Bellman optimality equation:
 […]
@@
 $\mathbb{E}\Bigg{[}\sum_{k=0}^{F_{i+1}-1}\bigg{(}\varepsilon\big{(}b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big{)}-\varepsilon_{\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$
-$+\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]}\big{\}}\big{\}}.$ (17)
+$+\mathbb{E}\big{[}h_{2}\big{(}b_{i+1}+T_{i+1}(l_{i+1})+F_{i+1},l_{i+1},C_{i+1}\big{)}\big{|}C_{i}=c_{i}\big{]}\Bigg{\}}\Bigg{\}}.$ (17)
 Proof Sketch.
-Theorem 2 is proved motivated by the observation that the optimal waiting time $\tau_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ for any state $\big{(}\delta,d,c_{i}\big{)}$ can be determined by solving separate optimization problems, given the optimal packet length $l_{i+1}^{*}$.
-Subsequently, the solution to the waiting time problem is derived as an index-based threshold rule, which allows us to simplify the Bellman optimality equation (IV-B) to (2).
-The detailed proof is provided in Appendix E.
+Theorem 2 is proved based on the observation that the optimal waiting time $\tau_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ for any state $\big{(}\delta,d,c_{i}\big{)}$ can be determined by solving separate optimization problems, provided that the optimal packet length $l_{i+1}^{*}$ is known.
+We derive a solution for the separate waiting time optimization problem as an index-based threshold rule, which, in turn, enables the simplification of the Bellman optimality equation (IV-B) to (2).
+The detailed proof is presented in Appendix E.
 ∎
-Both Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [20, Chapter 11.4.4].
-Policy iteration is efficient for solving SMDPs [17, 20].
-The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B).
-The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2).
+The Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [23, Chapter 11.4.4].
+In particular, policy iteration is efficient for solving SMDPs [17, 23].
+The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B): The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2).
 The policy evaluation step is identical for both equations.
 V Simulations
-To evaluate the performance of the optimal scheduling policies outlined in Theorems 1 and 2, we consider two experiments: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which provides a trace-driven evaluation.
+To evaluate the performance of the optimal scheduling policies in Theorems 1 and 2, we consider two experiments: (i) remote inference of auto-regressive (AR) processes, which offers a model-based evaluation; and (ii) cart-pole state prediction, which provides a trace-driven evaluation.
 V-A Model-Based Evaluation
-Control systems with memory and delay [43, 44, 45] and wireless channels [46] are often modeled as higher-order autoregressive (AR) linear time-invariant systems.
-Motivated by this, we consider a remote inference problem for an AR process, where the target evolves according to:
+Figure 3: Inference error of the AR process for AoI values ${\delta=1,2,\ldots,50}$ and packet lengths $l=1,2,4,6,8,10$.
+Higher-order auto-regressive (AR) linear time-invariant systems are widely used models for various practical scenarios, such as control systems incorporating memory and delay [43, 44, 45], and wireless communication channels [46].
+Motivated by this, we consider a remote inference problem for an AR process, where the target signal evolves according to:
 $Y_{t}=a_{1}Y_{t-1}+a_{2}Y_{t-2}+\ldots+a_{p}Y_{t-p}+W_{t},$ (18)
@@
 We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$.
-The remaining coefficients are zero.
+The remaining coefficients of the AR($10$) process are zero.
 We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$.
 In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$.
-Since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic, the optimal inference error is achieved by a linear MMSE estimator.
+A linear MMSE estimator achieves the optimal inference error since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic.
 Fig.
-3 shows the average inference error for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
-Figure 3: Inference error of the AR process experiment for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
-Using this AR process experiment, we illustrate the performance of the optimal policy derived for the scheduling problem (6)-(7) under time-invariant packet length selection.
-Let the number of delay states $N$ be $2$.
-The transmission delay $T_{i}(l)$ corresponding to each state is given by
-$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (19)
+3 shows the inference error for AoI values ${\delta=1,2,\ldots,50}$ and packet lengths $l=1,2,4,6,8,10$.
+Using this AR process, we illustrate the performance of the optimal policy derived for the scheduling problem specified in (6)-(7) under time-invariant packet length selection.
+Let the number of delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$.
+During the simulations, we assume $p_{12}=p_{21}$, ensuring that each delay state has a steady-state probability of $\frac{1}{2}$.
+The transmission delay $T_{i}(l)$ for each delay state is given by
+$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (19)
 where $\sigma$ is a scaling parameter.
-The feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$.
-We adjust the delay memory using the parameter $\alpha=p_{12}+p_{21}$, which is equal to the sum of transition probabilities of the Markov chain according to which the delay state evolves.
-We set $p_{12}=p_{21}$, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in the steady state for any $\alpha$.
-The maximum allowed packet length is 10, and the buffer size $B=75$.
+Additionally, the feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$.
+The maximum allowed packet length is $10$, and the buffer size is $B=75$.
+We control the delay memory by varying the sum of transition probabilities $\alpha=p_{12}+p_{21}$.
+When $\alpha=1$, i.e., ${p_{12}=p_{21}=\frac{1}{2}}$, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the $(i-1)$-th epoch.
+Consequently, the delay distribution is i.i.d..
+However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the $(i-1)$-th epoch becomes more informative about the delay state in the $i$-th epoch.
 Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $2$ with $\alpha=\frac{1}{20}$:
@@
   • 2.
-    Policy outlined in Theorem 1 for $l=1$.
+    Policy outlined in Theorem 1 with $l=1$.
   • 3.
     Generate-at-will, zero-wait, $l=1$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(1,1,\ldots)$.
-The result in Fig.
-4 highlights that the first policy achieves a time-average inference error up to 6 times smaller compared to the third policy.
-Although part of this gain can be obtained with the packet length $l=1$ by optimally determining the waiting times and buffer positions as in the second policy, a larger portion of the gain comes from selecting the appropriate packet length, accounting for the interplay between the transmission delay statistics and the packet length $l$.
+The result presented in Fig.
+4 highlights that the optimal policy achieves a time-average inference error up to six-times lower than that of the generate-at-will, zero-wait, $l=1$ policy.
+Part of this improvement can be attained using a packet length of $l=1$ with the waiting times and buffer positions determined as described in Theorem 1.
+However, a substantial portion of the performance gain comes from selecting the appropriate packet length, accounting for the interplay between the transmission delay statistics and the packet length $l$.
 Figure 4: Time-average inference error vs. the scaling parameter $\sigma$ of transmission delay $T_{i}(l)$.
@@
   • 1.
-    Policy outlined in Theorem 1 for $l=5$.
+    Policy outlined in Theorem 1 with $l=5$. This policy accounts for the delay memory.
   • 2.
-    Policy outlined in [17, Theorem 1] for a fixed packet length of $l=5$.
-During the simulations, this policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
+    Policy outlined in [17, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
 The result in Fig.
-(5) emphasizes the importance of considering delay memory when designing scheduling algorithms.
-It demonstrates that the first policy offers minimal improvement compared to the second policy when $\alpha$ is close to $1$.
-This is because the delay distribution becomes exactly IID at $\alpha=1$.
-However, as $\alpha$ deviates from this value, the delay memory increases, and the advantage of the first policy becomes apparent, with a significant performance gain of up to $11.6\%$.
+(5) emphasizes the importance of accounting for delay memory when designing scheduling algorithms.
+The performance of both policies is quite similar when the delay state of the network is almost i.i.d., i.e., when $\alpha$ is around $1$.
+However, as $\alpha$ deviates from this value, the delay memory increases, and the advantage of the policy in Theorem 1 becomes apparent, with a significant performance gain of up to $11.6\%$.
 Figure 5: Time-average inference error vs. the sum of transition probabilities $\alpha=p_{12}+p_{21}$.
 V-B Trace-Driven Evaluation
+Figure 6: Structure of the cart-pole experiment.
+Ensuring the safe operation of robotic systems often requires evaluating their state based on observations from remote sources.
+This makes the prediction problem critical for monitoring the safety of such systems.
+We consider a specific scenario, illustrated in Fig.
+6, where a camera remotely monitors a control system, e.g., cart-pole.
+The camera captures video frames and transmits them to a receiver over a network with distinct delay states.
+Due to the high dimensionality of the video data, transmitting the sequence of frames takes multiple time slots.
+As a result, the most recently delivered sequence of video frames is generated $\delta$ time slots ago.
+Upon reception, the frames undergo data pre-processing to extract the cart position and the pole angle information, ${X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta-l+1},Z_{t-\delta-l+1})\big{)}}$.
+Subsequently, a predictor uses the most recent available information to estimate the current pole angle and cart position ${Y_{t}=(\psi_{t},Z_{t})}$.
+Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths ${l=1,2,4,6,8,10}$.
 We consider the OpenAI CartPole-v1 task [47], where a Deep Q-Network (DQN) reinforcement learning algorithm [48] is used to control the force applied to a cart to prevent the attached pole from falling over.
@@
 This dataset was generated by simulating 10,000 episodes of the OpenAI CartPole-v1 environment using the DQN controller.
-The objective is to predict the pole angle and position $Y_{t}=(\psi_{t},Z_{t})$ at time $t$ based on the feature vector $X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta+l-1},Z_{t-\delta+l-1})\big{)}$, which represents a sequence of past pole angles and cart positions generated $\delta$ time slots ago, with a sequence length of $l$.
-Figure 6: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$.
-Ensuring the safe operation of industrial and robotic systems often requires evaluating their state based on observations from remote sources.
-Consequently, the prediction problem is critical for monitoring the safety of these systems.
-We focus on a specific scenario where a camera remotely monitors a cart-pole system.
-The camera captures video frames and transmits them to a receiver.
-Due to the high dimensionality of the video data, transmitting the sequence of frames requires multiple time slots.
-Consequently, the most recently delivered sequence of video frames is generated $\delta$ time slots ago.
-Once received, the frames undergo signal processing to extract relevant information about the cart’s position and the pole’s angle.
-We assume this signal processing is accurate.
-Then, a predictor takes the most recent available observations and predicts the current pole angle and cart position $Y_{t}=(\psi_{t},Z_{t})$.
 The predictor in this experiment is a Long Short-Term Memory (LSTM) neural network consisting of one input layer, one hidden layer with 64 LSTM cells, and a fully connected output layer.
@@
 The inference error is shown in Fig.
-6 for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$.
-We observe that the inference error is not a monotonic function of AoI $\delta$ for a given feature vector length $l$.
-Additionally, the inference error decreases as the feature vector length $l$ increases for a fixed AoI value $\delta$.
-Using this cart-pole experiment, we illustrate the performance of the optimal policy outlined in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection.
-We again consider two delay states, i.e., ${N=2}$.
-The transmission delay $T_{i}(l)$ is given by
-$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (20)
-where $\sigma$ is a scaling parameter.
-The feedback delay $F_{i}$ is identical to that in the AR process experiment.
-We set the parameter $\alpha=p_{12}+p_{21}=\frac{1}{20}$.
-The transition probabilities $p_{12}$ and $p_{21}$ are equal, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in steady state for any $\alpha$.
+7 for AoI values $\delta=1,2,\ldots,80$ and packet lengths $l=1,2,4,6,8,10$.
+We observe that the inference error is not a monotonic function of AoI $\delta$ for a given packet length $l$.
+Moreover, the inference error decreases as the packet length $l$ increases for a fixed AoI value $\delta$.
+Using this cart-pole experiment, we illustrate the performance of the optimal policy in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection.
+We consider two delay states $N=2$.
+The delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$.
+We set $p_{12}=p_{21}$ to ensure equal steady-state probabilities for both delay states.
+The transmission delay $T_{i}(l)$ for each delay state is given by
+$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (20)
+where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the sum of transition probabilities as $\alpha=p_{12}+p_{21}=\frac{1}{20}$.
 (a) $\sigma=\frac{1}{2}$
@@
 Fig.
-7(a) and Fig.
-7(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$, respectively.
+8(a) and Fig.
+8(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$, respectively.
 As the buffer size increases, the performance improvement diminishes and becomes less significant.
@@
 VI Conclusion
-This paper studies a goal-oriented communication design problem for remote inference, where an intelligent model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location.
+This paper studies a goal-oriented communication design problem for remote inference, where an intelligent neural network model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location.
 We derive two optimal scheduling policies under time-invariant and time-variable packet length selection.
 These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory.
-Finally, through the simulations, we demonstrate the significant benefits of adopting such a goal-oriented communication strategy for remote inference.
+Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
 Acknowledgment
@@
 $\mathbb{E}\Bigg{[}\sum_{k=0}^{T_{i+1}(l)-1}\bigg{(}\varepsilon\big{(}\delta+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}.$ (22)
-By some algebraic manipulations, we can obtain
+By using the linearity of expectation, we can obtain
 $\min_{\tau_{i+1}\in\{1,2,\ldots\}}\mathbb{E}\Bigg{[}\sum_{k=0}^{\tau_{i+1}-1}\bigg{(}\varepsilon\big{(}\delta+T_{i+1}(l)+k,l\big{)}-\varepsilon_{l,\text{opt}}\bigg{)}\Bigg{|}C_{i}=c_{i}\Bigg{]}$

### v4

@@
 Abstract
-We study a setting where an intelligent model (e.g., a pre-trained neural network) predicts the real-time value of a target signal using data samples transmitted from a remote source according to a scheduling policy.
-The scheduler decides on i) the age of the samples to be sent, ii) when to send them, and iii) the length of each packet (i.e., the number of samples contained in each packet).
-The dependence of inference quality on the Age of Information (AoI) for a given packet length is modeled by a general relationship.
-Previous work assumed i.i.d.
+We study a setting where an intelligent model (e.g., a pre-trained neural network) infers the real-time value of a target signal using data samples transmitted from a remote source.
+The transmission scheduler decides (i) the freshness of packets, (ii) their length (i.e., the number of samples they contain), and (iii) when they should be transmitted.
+The freshness is quantified using the Age of Information (AoI), and the inference quality for a given packet length is a general function of AoI.
+Previous works assumed i.i.d.
 transmission delays with immediate feedback or were restricted to the case where inference performance degrades as the input data ages.
@@
 We model this as an infinite-horizon average-cost Semi-Markov Decision Process.
-We obtain a closed-form solution that decides on (i) and (ii) for any constant packet length.
-The solution for when to send is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver.
-The age of the packet selected is a function of the delay state.
-We separately optimize the value of the constant length.
-We also develop an index-based threshold policy for the variable length case, which allows a complexity reduction.
-In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
-Index Terms:
-Remote Inference, Goal-oriented Communication, Delay with Memory, Markovian Delay, Two-way Delay, Age of Information, and Dynamic Programming.
+We obtain a closed-form solution that decides on (i) and (iii) for any constant packet length.
+The solution for when to transmit is an index-based threshold policy, where the index function is expressed in terms of the delay state and AoI at the receiver.
+In contrast, the freshness of the selected packet is a function of only the delay state.
+We then separately optimize the value of the constant packet length.
+Moreover, we also develop an index-based threshold policy for the time-variable packet length case, which allows a complexity reduction.
+In simulation results, we observe that our goal-oriented scheduler drops inference error down to one-sixth with respect to the age-based scheduling of unit-length packets.
 I Introduction
@@
 These models have played a key role in enabling certain technologies as digital twins, industrial robotics, and self-driving cars.
-However, these models rely on data to be delivered to them, which in the remote computation context, depends on the network delivering the relevant data in a timely fashion.
-This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2].
+However, these models rely on data to be delivered to them, which, in the remote computation context, depends on the network delivering the relevant data in a timely fashion.
+This is especially interesting in the case of unavoidable and time-varying transmission delays and is within the realm of Goal-oriented Networking for AI [2, 3].
 An extreme example of monitoring over long-distance network connections is a spacecraft digital twin.
@@
 The digital twin can be monitored over the Internet, and device settings in the physical experiment configuration on site can be remotely changed on demand directly on the twin.
-We focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [5, 6], which are often essential for such applications.
-Goal-oriented communication has been recently introduced to address the scalability problem in next-generation communication networks.
-This concept emphasizes that networks can facilitate the effective accomplishment of the task at the destination rather than merely addressing the transmission problem [7], which aims to reliably transmit data produced by a source.
-Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that data samples most significant to the computation at the destination are delivered in a timely manner [8].
+In this paper, we focus on goal-oriented communication design for real-time estimation and inference performed by intelligent models [6, 7, 8], which are often essential for such applications.
+Goal-oriented communication has been recently introduced to address the scalability problem in next-generation massive machine-type communication networks.
+This concept emphasizes that networks can facilitate the effective accomplishment of tasks at the destination rather than merely addressing the transmission problem [9], which aims to reliably transmit data produced by a source.
+Solving the effective communication problem with efficient use of network resources requires combining the data generation and transmission processes, ensuring that the samples most relevant to the computation at the destination are delivered in a timely manner [2].
 This implies that the communication link should select which samples to transmit based on the state of the network (e.g., the delay state).
@@
 This function is all that is required for the link layer to operate in a goal-oriented manner.
-This work builds upon the findings presented in a series of papers [5, 16, 6, 17, 1, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
-In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$, based on a data packet, $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$, transmitted from a remote location.
+This work builds upon the findings presented in a series of papers [6, 17, 7, 8, 18, 19, 20], which first demonstrated the effectiveness of AoI function as a surrogate metric for enhancing timely communication for remote estimation and inference.
+In this paper, we study a remote inference problem, where an intelligent, pre-trained model at the receiver is utilized to infer the real-time value of a target signal, $Y_{t}$.
+This inference is based on the most recently received data packet, denoted as $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$}, which has been transmitted from a remote location.
 The data packet $X_{t-\delta}^{l}$ comprises samples of the source signal $V_{t}$ at the transmitter.
@@
 For instance, $\Upsilon$ may represent a combination of the communication delay between the controller and the actuator and the actuation delay in a physical networked control system.
-An example of this scenario is illustrated in the leader-follower robot setup in [6, Fig.
+An example of this scenario is illustrated in the leader-follower robot setup in [7, Fig.
 2].
-Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [6, Fig.
+Additionally, non-monotonicity also occurs when the relationship between $V_{t}$ and $Y_{t}$ exhibits periodicity, as shown in the temperature prediction example in [7, Fig.
 4].
 In such cases, when communication delays prevent the transmission of sufficiently fresh samples to the receiver, transmitting aged samples can exploit the periodicity to enhance performance.
-Moreover, the inference error is non-increasing with packet length for a given AoI [17], as additional information can only improve the estimation accuracy.
+The inference error is non-increasing with packet length for a given AoI [8], as additional information can only improve the estimation accuracy.
 However, this improvement comes at the cost of longer transmission delays.
-In our communication design, we investigate the interplay between packet length and transmission delay, as well as the possibly non-monotonic relationship between inference error and AoI.
+In our communication design, we investigate the interplay between packet length and transmission delay, along with the possibly non-monotonic relationship between inference error and AoI.
 Finally, next-generation communication networks are expected to accommodate numerous connections, resulting in multiple routes between any pair of nodes.
@@
 We consider both time-invariant and time-variable packet length selection to address the varying computational capabilities of practical systems.
-The derived optimal scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
-  • We model the learning and communication co-design problem under time-invariant packet length selection as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP).
+The derived optimal packet selection and scheduling policies minimize the time-average inference error for a given, possibly non-monotonic, inference error function corresponding to a particular remote inference application, thereby making the communication goal-oriented.
+  • We formulate the learning and communication co-design problem under time-invariant packet length selection as a two-layer nested optimization problem.
+The inner layer determines the freshness and transmission times of data packets for a given constant packet length $l$ and is modeled as an infinite-horizon average-cost Semi-Markov Decision Process (SMDP).
 Such problems are often solved using dynamic programming [23, 24] and typically do not have closed-form solutions.
-However, we derive a closed-form solution to this problem for any constant packet length, presented in Theorem 1.
-The solution consists of two main parts: (i) the AoI of data packets to be sent are determined as a stationary function of the network state, i.e., the delay state, and (ii) the transmission instants are governed by an index-based threshold policy.
-The index function is expressed in terms of the delay state of the network and the AoI at the receiver.
-We then separately optimize the value of constant packet length.
-  • We also model the learning and communication co-design problem under time-variable packet length selection as an infinite-horizon average-cost SMDP.
+However, we derive a closed-form solution to this inner layer problem, presented in Theorem 1.
+The solution consists of two main parts: (i) the freshness of the data packets to be transmitted is determined by a stationary function of the network state, i.e., the delay state, in the previous epoch, and (ii) the transmission instants are controlled by an index-based threshold policy.
+The index function is expressed in terms of the delay state of the network in the previous epoch and the AoI at the receiver.
+We then separately optimize the value of the constant packet length $l$ in the outer layer.
+  • We model the learning and communication co-design problem under time-variable packet length selection directly as a single infinite-horizon average-cost SMDP.
 We then formulate a Bellman optimality equation and derive a structural result regarding the optimal solution.
@@
   • We conduct two experiments to evaluate the performance of our optimal scheduling policies: (i) remote inference of auto-regressive (AR) processes, offering a model-based evaluation; and (ii) cart-pole state prediction, providing a trace-driven evaluation.
-In simulation results, we observe that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
+In simulation results, we observe that our goal-oriented scheduler drops inference error down to one-sixth with respect to the age-based scheduling of unit-length packets.
 I-A Related Work
@@
 Recent research endeavors have revealed that the performance of real-time applications can be modeled as non-linear functions of AoI, leading to the study of optimizing these non-linear functions in control system scenarios [28, 36], remote estimation [17, 26, 31], and remote inference [22, 6, 8, 7].
-While a number of studies have analyzed AoI in queuing models, closest to the spirit of this paper is the control of AoI via replacement of exogenous data arrivals with the generation of data ”at will” [37, 18, 38, 39, 19, 27, 28, 29].
+While many studies have analyzed AoI in queuing models, closest to the spirit of this paper is the control of AoI via replacement of exogenous data arrivals with the generation of data ”at will” [37, 18, 38, 39, 19, 27, 28, 29].
 A generalization of this approach is to incorporate jointly optimal sampling and scheduling policies to control not only AoI but a more sophisticated end-to-end distortion criterion by using AoI as an auxiliary parameter [17, 26, 33, 8, 6, 7].
@@
 This paper extends the problem formulations in [7, 8] to more practical scenarios and develops a goal-oriented communication strategy that minimizes the average inference error for remote inference under two-way delay (i.e., incorporating random feedback delay [40, 41]) that varies significantly with memory.
+The earlier version of this paper [1] is a special case of the learning and communication co-design problem under time-invariant packet length selection discussed in Section III.
+In [1], the scheduler is restricted to transmitting a single sample at each transmission, i.e., the packet length is fixed to $l=1$.
+In contrast, the current paper allows the scheduler to form packets comprising multiple consecutive samples for transmission.
+Packets containing a larger number of samples can yield better inference performance.
+However, the associated increase in transmission delay makes such packets staler upon arrival, which may eventually degrade inference performance.
+As a result, this paper captures the trade-off between the volume of data transmitted and the resulting increase in transmission delay while optimizing the task performance at the destination.
+Furthermore, [1] established the optimal threshold-based waiting-time rule for the time-invariant packet length $l=1$ and showed that the freshness of the data packets to be transmitted is a function only of the network delay state in the previous epoch.
+The present paper extends these results by generalizing the waiting-time rule to arbitrary time-invariant packet lengths $l$ (not limited to $l=1$) and by deriving an exact closed-form expression for the freshness of the data packets to be transmitted.
+Finally, this paper also investigates scenarios where the packet length can be adjusted over time based on the state of the scheduler in Section IV.
 II System Model
@@
 The samples are stored in a buffer containing the most recent $B$ samples $\big(V_{t},V_{t-1},\ldots,V_{t-B+1}\big)$, waiting to be sent to the receiver.
-On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of a target signal $Y_{t}$ of interest.
+On the receiver side, a predictor, e.g., a pre-trained neural network, is used to infer the real-time value of the target signal $Y_{t}$ of interest.
 In practical time-series forecasting algorithms, the predictor needs to use multiple consecutive samples, e.g., $X_{t-b}^{l}=(V_{t-b},V_{t-b-1},\ldots,V_{t-b-l+1})$, to infer $Y_{t}$.
@@
 This medium access model is called “selection-from-buffer,” which was proposed recently in [6, 7].
-Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
+Such remote inference systems are crucial for numerous real-time applications, including sensor networks, airplane/vehicular control, robotics networks, and digital twins, where timely and accurate remote monitoring and control are essential for maintaining system stability and performance.
 We assume that the system starts operation at time slot $t=0$.
@@
 At time slot $t=S_{i}$, the scheduler selects a group of consecutive samples from the buffer and forms the packet $X_{S_{i}-b_{i}}^{l_{i}}=\big(V_{S_{i}-b_{i}},V_{S_{i}-b_{i}-1},\ldots,V_{S_{i}-b_{i}-l_{i}+1}\big)\in\mathcal{V}^{l_{i}}$.
-Here, $l_{i}\in\{1,\ldots,B\}$ represents the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$ while $b_{i}$ represents the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer.
+Here, $l_{i}\in\{1,\ldots,B\}$ specifies the sequence length of the $i$-th packet $X_{S_{i}-b_{i}}^{l_{i}}$, while $b_{i}$ denotes the relative position of the sample $V_{S_{i}-b_{i}}$ in the buffer.
+The latter parameter $b_{i}$ controls the freshness of the samples included in the packet.
 When $B=1$, the selection-from-buffer model reduces to the generative-at-will model which has been studied extensively in previous works [37, 18, 38, 39, 19, 27, 28, 29].
@@
 Upon delivery, the receiver sends an acknowledgment (ACK) back to the transmitter, which is received at time slot $A_{i}$.
-We assume that the transmitter always waits for feedback before submitting a new packet.
-In other words, the transmitter remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$.
+We assume that the scheduler always waits for feedback before transmitting a new data packet.
+In other words, the scheduler remains silent during the time slots between $D_{i}$ and $A_{i}$ such that $S_{i}<D_{i}<A_{i}\leq S_{i+1}$ for all $i$.
+This assumption clearly restricts the throughput, the maximization of which has been the main objective in conventional user-centric, pipelined communication systems.
+However, under the concept of goal-oriented communication, which has been introduced to address the scalability problem in massive machine-type networks, the main objective is not throughput maximization but ensuring the effective execution of machine tasks at the destination while operating under limited communication resources.
+From this perspective, allowing multiple packets to be simultaneously in flight may increase network load without yielding proportional improvements in task performance, particularly in dense environments where congestion and contention are the primary concerns.
+Consequently, the one-packet-in-flight (stop-and-wait) assumption serves as a scenario-driven modeling choice that captures conservative communication behavior suitable for the class of systems considered in this work.
+II-A Communication Network Model
+Next-generation communication networks are expected to accommodate multiple routes between a transmitter–receiver pair through a variety of intermediate network components, such as LEO satellites, inter-satellite links, and ground gateways.
+The experienced delay may therefore vary significantly depending on the specific components traversed by a packet and the manner in which it is handled during transmission.
+For example, when a packet is delivered entirely via inter-satellite links under moderate network load, the dominant delay component is propagation delay.
+In contrast, when inter-satellite link unavailability requires routing through intermediate ground stations, or when packets undergo store-and-forward transmission over long distances in LEO satellites under heavy network load, the resulting delay may involve either substantially larger propagation components or considerable store-and-forward waiting times.
+In such networks, the receiver can characterize the realized end-to-end delay behavior through a detection rule that is consistent with its observation capabilities (e.g., based on measured packet arrival times).
+The precision of this characterization depends on the receiver’s observation capability.
+Furthermore, as in any practical communication system, physical-layer transmission errors and link-layer retransmissions may occur along these routes.
+In this work, we adopt a high-level network-layer perspective, where such effects are incorporated into the end-to-end transmission delay.
+Specifically, the delay experienced by a packet accounts for processing and propagation delays, as well as potential local retransmissions and error recovery mechanisms within the network.
+This modeling choice allows us to focus on the scheduler’s decision-making at the source, rather than on lower-layer protocol operations.
+Figure 2: An illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
+Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$.
+Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay.
+In addition, let $C_{i}$ denote the network delay state in the $i$-th epoch, with its realization $c_{i}$.
 Fig.
 2 provides an illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
-II-A Communication Network Model
-Figure 2: An illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
-Consider that the communication network may have multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver.
-The network directs each submitted packet to one of these available routes.
-If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route.
-Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$.
-Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay.
-Additionally, let $C_{i}$ represent the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization.
-The delay state $C_{i}$ is determined by the route used in the $i$-th packet transmission epoch.
-Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length.
+The network delay state $C_{i}$ in the $i$-th epoch can take values in $\{1,2,\ldots,N\}$, each representing a distinct delay regime that the network may experience during the transmission of the $i$-th data packet and that can be distinguished by the receiver.
+Each delay regime is associated with different transmission and feedback delay distributions, and the transmission delay also varies with the packet length.
 The delay is at least one time slot for each packet or ACK transmission.
-We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$.
+We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$ and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$.
 The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise.
-The receiver detects the delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message.
+The receiver detects the network delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message.
 The ACK notifies the transmitter of (i) the delay state experienced during the previous packet transmission, (ii) the completion of the previous packet transmission, and (iii) that the next packet can be submitted to the network.
@@
 The AoI $\delta$ denotes the time difference between the generation of $\tilde{Y}_{0}$ and $\tilde{X}_{-\delta}^{l}$.
-These trained neural networks are used to predict the real-time value of the target signal $Y_{t}$ on the receiver side.
-We assume that the process ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ is stationary, and that the processes ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ and ${\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}}$ are independent.
-The average inference error at time slot $t$, given $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
+These pre-trained neural networks are used at the destination to predict the real-time value of the target signal $Y_{t}$ based on the most recently received data packet.
+We assume that the stochastic process $\{\big(Y_{t},V_{t}\big),t=0,1,\ldots\}$} is stationary, implying that the statistical relationship between the source signal and the target signal is time-invariant.
+In addition, we further assume that the process $\{\big(Y_{t},V_{t}\big),t=0,1,\ldots\}$} is independent of the process $\{\big(\Delta(t),d(t)\big),t=0,1,\ldots\}$}.
+Accordingly, the average inference error at time slot $t$, conditioned on $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
 $\varepsilon\big(\delta,l\big)=\mathbb{E}_{Y,X^{l}\sim P_{Y_{t},X_{t-\delta}^{l}}}\big[L\big(Y,\phi^{*}_{l}\big(\delta,X^{l}\big)\big)\big],$ (4)
@@
 III-A Co-design Problem Formulation
-Let $\Pi_{l}$ represent the set of all causal scheduling policies $\pi=(f,\ell,g)$ with a time-invariant packet length sequence =${\ell=\{l,l,\ldots\}}$ for $l\in\{1,2,\ldots,B\}$:
+Let $\Pi_{l}$ represent the set of all causal scheduling policies $\pi=(f,\ell,g)$ with a time-invariant packet length sequence $\ell=\{l,l,\ldots\}$} for $l\in\{1,2,\ldots,B\}$:
 $\Pi_{l}=\{\pi\in\Pi:l=l_{2}=l_{3}=\ldots\}.$ (5)
@@
 Moreover, the outer problem (7) optimizes the packet length and involves a straightforward search over the integer values $\{1,2,\ldots,B\}$.
-Problem (6)-(7) is a more general formulation than that considered in [17, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in multi-route communication networks.
+Problem (6)-(7) is a more general formulation than that considered in [8, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in next-generation communication networks.
 III-B Optimal Solution to (6)-(7)
@@
 Step 2: Since the delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining $b_{i+1}^{*}$ for all $i$, we can construct a new SMDP for determining $b_{i+1}^{*}$ using $c_{i}$ as the state.
-In the new SMDP, we fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$.
+In the new SMDP, we fix the waiting time decisions by using the optimal threshold rule $\tau_{l}\big(\delta,c_{i}\big)$.
 Moreover, in the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$.
@@
 The index function $\gamma\big(\cdot\big)$, defined in equation (9), can be readily computed for any state $\big(\delta,c_{i}\big)$ using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$.
-Although the index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ closely resembles the waiting time rule provided in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$.
+Although the index-based threshold rule $\tau_{l}\big(\delta,c_{i}\big)$ closely resembles the waiting time rule provided in [8, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$.
 Because problem (6)-(7) involves non-zero feedback delay and time-varying delay distribution, the existence of such a simple waiting time rule was not evident and required additional technical efforts for its derivation.
@@
 For the case of i.i.d.
-transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [17, Theorem 1].
+transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [8, Theorem 1].
 However, when the feedback delay is non-zero and the delay state $C_{i}$ changes according to a Markov chain, achieving optimal performance requires that the optimal buffer position $b_{i+1}^{*}$ is a function of the delay state $C_{i}=c_{i}$, and this function is exactly given by (1).
@@
 Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the following simplified Bellman optimality equation:
-[…]
+$h_{2}\big(\delta,d,c_{i}\big)=\min_{\begin{subarray}{c}l_{i+1}\in\{1,2,\ldots,B\}\end{subarray}}\Bigg\{$
 $\mathbb{E}\Bigg[\sum_{k=0}^{\tau\big(\delta,d,l_{i+1},c_{i}\big)+T_{i+1}(l_{i+1})-1}\bigg(\varepsilon\big(\delta+k,d\big)-\varepsilon_{\text{opt}}\bigg)\Bigg|C_{i}=c_{i}\Bigg]$
-[…]
+$+\min_{b_{i+1}\in\{0,1,\ldots,B-l_{i+1}\}}\Bigg\{$
 $\mathbb{E}\Bigg[\sum_{k=0}^{F_{i+1}-1}\bigg(\varepsilon\big(b_{i+1}+T_{i+1}(l_{i+1})+k,l_{i+1}\big)-\varepsilon_{\text{opt}}\bigg)\Bigg|C_{i}=c_{i}\Bigg]$
@@
 In particular, policy iteration is efficient for solving SMDPs [23, 8].
-The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B): The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2).
+The time complexity for solving the simplified Bellman equation (2) via policy iteration is significantly lower than that for solving (IV-B).
+Specifically, the policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2).
 The policy evaluation step is identical for both equations.
@@
 Higher-order auto-regressive (AR) linear time-invariant systems are widely used models for various practical scenarios, such as control systems incorporating memory and delay [43, 44, 45], and wireless communication channels [46].
-Motivated by this, we consider a remote inference problem for an AR process, where the target signal evolves according to:
+Motivated by this, we consider a remote inference problem for an AR process, where the target signal $Y_{t}\in\mathbb{R}$ evolves according to:
 $Y_{t}=a_{1}Y_{t-1}+a_{2}Y_{t-2}+\ldots+a_{p}Y_{t-p}+W_{t},$ (18)
-where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$ and $Y_{t}\in\mathbb{R}$.
-Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of the target $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$.
-We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$.
+where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$.
+Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$.
+In this experiment, we set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$.
 The remaining coefficients of the AR($10$) process are zero.
 We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$}.
-In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$.
+The goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$.
 A linear MMSE estimator achieves the optimal inference error since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic.
@@
 Using this AR process, we illustrate the performance of the optimal policy derived for the scheduling problem specified in (6)-(7) under time-invariant packet length selection.
-Let the number of delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$.
+For this experiment, we set the maximum allowable packet length to $l=10$ and the buffer size to $B=75$.
+Let the number of network delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $\{p_{jk}\}_{j,k\in\{1,2\}}$.
 During the simulations, we assume $p_{12}=p_{21}$, ensuring that each delay state has a steady-state probability of $\frac{1}{2}$.
-The transmission delay $T_{i}(l)$ for each delay state is given by
+We assume a simple transmission delay model that scales with the packet length and the current network conditions. The transmission delay $T_{i}(l)$ for each delay state is given by
 $T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (19)
-where $\sigma$ is a scaling parameter.
-Additionally, the feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$.
-The maximum allowed packet length is $10$, and the buffer size is $B=75$.
+where $\sigma$ is a scaling parameter. On the other hand, the corresponding feedback delay $F_{i}$ is given by
+$F_{i}=\begin{cases}1,&\text{if }c_{i}=1,\\ 3,&\text{if }c_{i}=2.\end{cases}$ (20)
 We control the delay memory by varying the sum of transition probabilities $\alpha=p_{12}+p_{21}$.
-When $\alpha=1$, i.e., ${p_{12}=p_{21}=\frac{1}{2}}$, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the $(i-1)$-th epoch.
+When $\alpha=1$, i.e., $p_{12}=p_{21}=\frac{1}{2}$}, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the (i1)(i-1)-th epoch.
 Consequently, the delay distribution is i.i.d..
-However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the $(i-1)$-th epoch becomes more informative about the delay state in the $i$-th epoch.
-Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $2$ with $\alpha=\frac{1}{20}$:
+However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the (i1)(i-1)-th epoch becomes more informative about the delay state in the $i$-th epoch.
+Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 0 to $2$ with $\alpha=\frac{1}{20}$:
   • 1.
@@
 Figure 4: Time-average inference error vs. the scaling parameter $\sigma$ of transmission delay $T_{i}(l)$.
-Fig. (5) presents the time-average inference error achieved by the following two scheduling policies as the parameter $\alpha$ varies from 00 to $2$ with $\sigma=\frac{5}{2}$:
+Fig. (5) presents the time-average inference error achieved by the following two scheduling policies as the parameter $\alpha$ varies from 0 to $2$ with $\sigma=\frac{5}{2}$:
   • 1.
@@
   • 2.
-    Policy outlined in [17, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
+    Policy outlined in [8, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
 The result in Fig.
@@
 We consider a specific scenario, illustrated in Fig.
-6, where a camera remotely monitors a control system, e.g., cart-pole.
+6, where a camera remotely monitors a control system, e.g., a cart-pole.
 The camera captures video frames and transmits them to a receiver over a network with distinct delay states.
@@
 Subsequently, a predictor uses the most recent available information to estimate the current pole angle and cart position $Y_{t}=(\psi_{t},Z_{t})$}.
-Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths ${l=1,2,4,6,8,10}$.
+Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths l=1,2,4,6,8,10{l=1,2,4,6,8,10}.
 We consider the OpenAI CartPole-v1 task [47], where a Deep Q-Network (DQN) reinforcement learning algorithm [48] is used to control the force applied to a cart to prevent the attached pole from falling over.
@@
 Using this cart-pole experiment, we illustrate the performance of the optimal policy in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection.
-We consider two delay states $N=2$.
-The delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$.
-We set $p_{12}=p_{21}$ to ensure equal steady-state probabilities for both delay states.
-The transmission delay $T_{i}(l)$ for each delay state is given by
-$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (20)
-where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the sum of transition probabilities as $\alpha=p_{12}+p_{21}=\frac{1}{20}$.
-(a) $\sigma=\frac{1}{2}$
-(b) $\sigma=\frac{5}{2}$
-Figure 8: Time-average inference error vs. the buffer size B for ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$.
+We consider a slightly more diverse network environment in this experiment by assuming three network delay states, i.e., $N=3$.
+Accordingly, the delay state $C_{i}$ in the $i$-th epoch follows a three-state Markov chain with transition probabilities $\{p_{jk}\}_{j,k\in\{1,2,3\}}$.
+We set the self-transition probabilities to $p_{11}=p_{22}=p_{33}=1-\alpha$, and all remaining transition probabilities to $\frac{\alpha}{2}$.
+Due to the symmetry of the Markov chain, the steady-state probability of each delay state is $\frac{1}{3}$.
+Throughout the simulation, we set $\alpha=\frac{1}{50}$.
+The transmission delay corresponding to each state has two components: (i) a constant $10$ time slot component, and (ii) a component that scales depending on the packet length and the network delay condition.
+For example, the first component may represent the unavoidable, relatively large propagation delay in non-terrestrial networks.
+The other part may represent the varying levels of store-and-forward operation of the satellites, as well as the increase in the transmission delay due to packet lengths.
+As a result, the transmission delay $T_{i}(l)$ for each delay state is given by
+$T_{i}(l)=10+\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ \lceil 2\sigma l\rceil,&\text{if }c_{i}=2,\\ \lceil 3\sigma l\rceil,&\text{if }c_{i}=3,\end{cases}$ (21)
+where $\sigma$ is a scaling parameter. On the other hand, the feedback delay $F_{i}$ for each delay state is given by
+$F_{i}=\begin{cases}1,&\text{if }c_{i}=1,\\ 2,&\text{if }c_{i}=2,\\ 3,&\text{if }c_{i}=3.\end{cases}$ (22)
+(a) $\sigma=\frac{1}{5}$
+(b) $\sigma=\frac{9}{5}$
+Figure 8: Time-average inference error vs. the buffer size B for $\sigma=\frac{1}{5}$} and $\sigma=\frac{9}{5}$}.
 Figure 9: Time-average inference error vs. the scaling parameter $\sigma$ of transmission delay $T_{i}(l)$.
@@
 8(a) and Fig.
-8(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$, respectively.
+8(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with $\sigma=\frac{1}{5}$} and $\sigma=\frac{9}{5}$}, respectively.
 As the buffer size increases, the performance improvement diminishes and becomes less significant.
 Therefore, we fix the buffer size $B$ at $10$ for this experiment.
-Fig. 9 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $3$:
+Fig. 9 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 0 to $2$:
   • 1.
@@
   • 2.
-    Generate-at-will, zero-wait, $l=1$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(1,1,\ldots)$.
+    Generate-at-will, zero-wait, $l=3$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(3,3,\ldots)$.
   • 3.
-    Generate-at-will, zero-wait, $l=5$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(5,5,\ldots)$.
+    Generate-at-will, zero-wait, $l=6$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(6,6,\ldots)$.
 The result in Fig.
 9 highlights the importance of adjusting the packet length over time in response to varying delay conditions.
-The optimal policy in Theorem 2 outperforms the other two policies, which use constant packet length, achieving up to $25\%$ reduction in average inference error.
+The optimal policy in Theorem 2 outperforms the other two policies, which use constant packet length, achieving up to $15.7\%$ reduction in average inference error.
+TABLE I: Normalized time complexity comparison for solving the original Bellman optimality equation (Orig.
+BE) (IV-B) and the simplified Bellman optimality equation (Simp.
+BE) (2), both with and without the pre-computed $\gamma(\cdot)$ in (9), using the policy iteration algorithm, for buffer sizes $B\in\{1,5,10\}$.
+Buffer size $B$
+Bellman equation $B=1$ $B=5$ $B=10$
+Orig. BE $\mathcal{T}_{1}$ $\mathcal{T}_{5}$ $\mathcal{T}_{10}$
+Simp. BE $\approx\mathcal{T}_{1}$ $0.8556\mathcal{T}_{5}$ $0.4527\mathcal{T}_{10}$
+Simp. BE + pre-computed $\gamma(\cdot)$ $\approx\mathcal{T}_{1}$ $0.8531\mathcal{T}_{5}$ $0.4464\mathcal{T}_{10}$
+We finally present an empirical comparison of the normalized time complexity for solving the original Bellman optimality equation (IV-B) and the simplified Bellman optimality equation (2), provided in Section IV, using the policy iteration algorithm.
+For the simplified Bellman optimality equation (2), the time complexity is reported for two cases: when the index function $\gamma(\cdot)$ in (9) is pre-computed and when it is not initially available.
+The results are summarized in Table I for buffer sizes $B\in\{1,5,10\}$.
+Let $\mathcal{T}_{k}$ denote the time required to solve the original Bellman equation for buffer size $B=k$.
+The corresponding solution times for the simplified Bellman equation, both with and without pre-computed $\gamma(\cdot)$, are expressed relative to $\mathcal{T}_{k}$.
+For a small buffer size of $B=1$, the time complexities of solving the two equations are nearly identical.
+This is because, when $B=1$, the action space is extremely limited: the only possible choice for the buffer position action $b_{i+1}$ at decision time $A_{i}$ is zero, and the only possible packet length action $l_{i+1}$ is one.
+Consequently, the reduction in complexity of the policy improvement step of the policy iteration algorithm is not observable.
+In contrast, as the buffer size increases, solving the simplified Bellman equation has notably lower time complexity than solving the original Bellman optimality equation, achieving more than a two-fold reduction at buffer size $B=10$.
+Moreover, the computation of the index function $\gamma(\cdot)$, which is required for solving the simplified Bellman equation, incurs only a small computational overhead.
 VI Conclusion
@@
 These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory.
-Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
+Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one-sixth with respect to age-based scheduling of unit-length packets.
 Acknowledgment
@@
 The AoI $\Delta(t)$ at time slot $t$ evolves according to equation (1).
-Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II).
+Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section II.
 The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@
 At each decision time $D_{i+1}$, the system state is represented only by the delay state $c_{i}$ in the $i$-th epoch.
-We fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$, and the action involves determining the buffer position $b_{i+1}$.
+We fix the waiting time decisions by using the optimal threshold rule $\tau_{l}\big(\delta,c_{i}\big)$, and the action involves determining the buffer position $b_{i+1}$.
 A detailed description of this SMDP can be found in Appendix D.
@@
 The AoI $\Delta(t)$ at time slot $t$ and the length $d(t)$ of the packet most recently delivered by time slot $t$ evolve according to equations (1) and (2), respectively.
-Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II).
+Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section II.
 The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@
 (ii) Action: At each decision time $D_{i+1}$, the scheduler determines the buffer position $b_{i+1}$ for the next packet to be transmitted.
-(iii) State and State Transitions: The system state at each decision time $D_{i+1}$ is the delay state $C_{i}=c_{i}$ in the $i$-th epoch, which evolves according to the finite-state ergodic Markov chain defined in Section (II).
+(iii) State and State Transitions: The system state at each decision time $D_{i+1}$ is the delay state $C_{i}=c_{i}$ in the $i$-th epoch, which evolves according to the finite-state ergodic Markov chain defined in Section II.
 The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.